### v1

# Unsupervised Multimodal Graph-based Model for Geo-social Analysis

Abstract
The systematic analysis of user-generated social media content, especially when enriched with geospatial context, plays a vital role in domains such as disaster management and public opinion monitoring. Although multimodal approaches have made significant progress, most existing models remain fragmented, processing each modality separately rather than integrating them into a unified end-to-end model. To address this, we propose an unsupervised, multimodal graph-based methodology that jointly embeds semantic and geographic information into a shared representation space. The proposed methodology comprises two architectural paradigms: a mono graph (MonoGrah) model that jointly encodes both modalities, and a multi graph (MultiGraph) model that separately models semantic and geographic relationships and subsequently integrates them through multi-head attention mechanisms. A composite loss, combining contrastive, coherence, and alignment objectives, guides the learning process to produce semantically coherent and spatially compact clusters. Experiments on four real-world disaster datasets demonstrate that our models consistently outperform existing baselines in topic quality, spatial coherence, and interpretability. Inherently domain-independent, the framework can be readily extended to diverse forms of multimodal data and a wide range of downstream analysis tasks.
keywords:
Machine learning, Artificial intelligence, Unsupervised machine learning, Natural language processing, Social media analysis, Disaster management
1 Introduction
Social media platforms continuously generate vast amounts of data, including posts and comments, likes, shares, images, and videos. This content is rich, diverse, and real-time, offering a valuable resource for understanding how people interact, express opinions, form communities, and respond to events. The widespread adoption of smartphones with built-in location-tracking capabilities has added another dimension to this landscape: geographic context. By enabling users to share their locations, these devices have facilitated the integration of geospatial data into online content, enhancing the contextual depth of social media analysis. A growing body of research underscores the importance of geo-location data in social media research. Prior studies have demonstrated its utility across a wide range of applications, from disaster response [goodchild2010crowdsourcing] to public health monitoring [Signorini2011]. In particular, geo-referenced content from micro-blogging platforms has been leveraged to model and monitor phenomena such as earthquakes [Havas2021Portability], floods [Huang2018Flood], refugee movements [Havas2021Refugee], and disease outbreaks [Stolerman2023Disease].
Previous social media research has often been focusing on semantic topic modeling, a fundamental analytical algorithmic technique to uncover latent thematic structures within large volumes of user-generated text. A range of scholarly works have utilized topic modeling to reveal latent thematic structures within such texts [Resch2018] [Hannigan2019]. However, the resulting topic representations are rarely contextualized with additional data modalities. Early efforts to analyze geo-social textual media in a multimodal setting typically employed sequential workflows, wherein each modality was processed in isolation through distinct modules [yin2011geographical].
While these methods have demonstrated utility in various applications, their modular and linear design often hinders the ability to capture complex interdependencies between modalities. Furthermore, this sequential structure can introduce dependencies between processing stages, rendering the overall output sensitive to the order in which analyses are performed. More recent approaches introduced advancements toward integrating multiple modalities such as time, geographic space, semantic topics, and sentiments in the analysis of social media [gonzalez2024neural] [hanny2025]. However, they still depend on numerous isolated components and fragmented intermediate processing steps, falling short of delivering a truly end-to-end processing methodology. Nonetheless, inherent heterogeneity of input modalities in this case has already added to the complexity of the task. For example, text is symbolic and sequential, while location data is numeric and geospatial, often resulting in misaligned feature representations and inconsistencies during fusion. The lack of a unified machine learning model constrains cross-modal interactions and introduces fragile interdependencies, whereby variations in preprocessing or processing order can significantly impact both performance and interpretability.
This paper presents a new methodology for multimodal geospatial machine learning, that jointly embeds semantic and geospatial information in a unified graph-based space, using end-to-end learning with a composite loss to generate clusters that are both semantically coherent and spatially interpretable. The proposed methodology is application-agnostic and can, without loss of generality, be adapted to diverse forms of multimodal data integration, making it well-suited for a wide range of data fusion scenarios and downstream tasks across domains. To capture cross-modal relationships in a holistic and structured manner, we construct a unified multimodal embedding space based on a graph structure, where we can adapt diverse data types, including text, geolocation, and optionally emotions or temporal signals, into a shared representation space. In this formulation, nodes in the graph represent user-generated post embeddings, while edges encode relational signals such as semantic similarity, geographic proximity, or user relations. This enables the model to attend to multimodal interactions in a context-aware fashion.
At the core of our methodology, we train an artificial neural network in an unsupervised manner to aggregate information from a node’s neighbors, where a modified message passing mechanism allows scaling by the corresponding edge weights. A key innovation of our model is the use of a composite loss function inspired by reference-less clustering evaluation approaches. To enable robust and discriminative representation learning, the loss integrates contrastive, coherence, and alignment objectives. The contrastive component preserves local semantic structure by pulling similar nodes closer to each other in the embedding space while pushing dissimilar ones apart. The coherence objective encourages the formation of well-defined clusters by maximizing intra-cluster similarity and inter-cluster separation. At the same time, the alignment term stabilizes training by ensuring each node remains close to the centroid of its assigned cluster, thus enhancing representational consistency. By embedding these objectives directly into the learning process, the model is able to discover clusters that are both semantically meaningful and spatially coherent. Our approach is specifically tailored to support disaster management efforts. In emergencies triggered by natural hazards such as floods, geo-social media provides rich multimodal data from affected communities, enhancing situational awareness for first responders. Through a case study of the 2017 Hurricane Harvey in the USA, we demonstrate how geographically fine-grained multimodal clusters of social media posts closely align with the disaster’s observed impacts, illustrating the practical value of our method in real-world crisis scenarios. The framework itself is designed to be application-agnostic and, without loss of generality, can be adapted to various forms of multimodal data integration. It is well-suited for diverse data fusion scenarios and downstream tasks across different domains. Empirical evaluations demonstrate that our approach outperforms existing baselines in topic coherence, spatial compactness, and interpretability, providing a scalable and principled framework for multimodal social media analysis.
In summary, this work makes the following distinct contributions:
  • Introducing a multimodal graph-based framework that seamlessly integrates multiple modalities (specifically semantic and geospatial data) into a unified representation space.
  • Proposing a domain-agnostic methodology for multimodal data analysis, underpinned by two unsupervised architectures, MonoGraph and MultiGraph, that adapt to diverse data characteristics and application needs, positioning the approach as a versatile framework for research and practice in fields such as disaster management, public health, and social media analytics.
  • Developing a composite loss function that combines contrastive, coherence, and alignment objectives, ensuring the generation of semantically coherent and spatially interpretable clusters.
  • Demonstrating the effectiveness of our approach through a real-world disaster case study, where geographically fine-grained multimodal clusters of social media posts closely align with observed impacts.
2 Related work
To contextualize our work, we first examine general topic modeling techniques and their geospatial adaptations, specifically tailored for analyzing social media data. We then delve into graph-based topic modeling approaches, emphasizing their strengths in capturing intricate relationships within the data.
2.1 Topic Modeling
Topic modeling is a widely used technique for discovering latent semantic topics within large document collections, serving as a variant of automated content analysis that assigns one or more “topic codes” to each document, thereby providing a structured thematic summary [hoyle2022content]. Among the foundational methods, Latent Dirichlet Allocation (LDA) [blei2003latent] introduced a fully generative probabilistic framework in which each document is modeled as a mixture over latent topics, and each topic is represented as a multinomial distribution over words. The generative idea of LDA has been extended to incorporate additional dimensions such as sentiment analysis [Lin2009Joint], temporal dynamics [Blei2006Dynamic], and social media user information [zhao2011comparing], enriching the interpretability and applicability of topic models across domains. However, LDA and its variants fundamentally rely on the bag-of-words assumption, which ignores word order and semantic relationships between words, limiting their effectiveness in handling short, noisy, or highly contextual texts such as tweets and social media posts [hong2010empirical]. To overcome these limitations, embedding-based topic modeling techniques have recently emerged. Dieng et al. [dieng2020topic] introduced the Embedded Topic Model (ETM), which integrates pre-trained word and sentence embeddings with LDA’s generative framework to better capture semantic relationships. Sia et al. [sia2020neural] further demonstrated that clustering high-dimensional semantic embedding vectors, derived from neural language models, can serve as a powerful alternative strategy, achieving performance comparable to traditional topic models. Building on these embedding-based approaches, Angelov [angelov2020top2vec] proposed Top2Vec, which leverages document embeddings and clustering algorithms to identify topics without requiring the number of topics to be pre-specified. This method utilizes techniques like Uniform Manifold Approximation and Projection (UMAP) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) for dimensionality reduction and flexible clustering. Grootendorst [grootendorst2022bertopic] further advanced this line of work with BERTopic, which incorporates transformer-based contextual embeddings (e.g. from BERT) combined with refined clustering strategies, consistently outperforming traditional LDA and Top2Vec in both coherence and interpretability. Despite their progress, these models generally follow a sequential and fragmented framework and often struggle to balance semantic accuracy, interpretability, and scalability, particularly when applied to the diverse and noisy nature of social media data.
2.2 Spatio-semantic Topic Modeling
As location-aware social media platforms proliferate, spatially grounded topic modeling has gained increasing attention. This line of work focuses on uncovering geographically coherent semantic patterns by integrating geospatial signals into the semantic topic modeling process. Early efforts extended classical models like LDA with spatial priors. For example, Spatial LDA (sLDA) [Yuan2013Who] and Geo-aware Topic Models [Hong2012Discovering] include location-based information as latent variables or observed metadata, enabling the discovery of region-specific topics. Similar models have been applied to a variety of domains, such as understanding natural disasters [Resch2018], flood mapping [Huang2018Flood], and epidemic monitoring [Stolerman2023Disease]. However, most spatial topic models treat location and text independently or as loosely coupled variables, which often leads to weak cross-modal alignment. In addition, spatial topic models are frequently built upon modular architectures, making them difficult to scale or generalize. More recent multimodal approaches attempt to combine geolocation, text, and other modalities, often through hybrid processing pipelines [gonzalez2024neural] [hanny2025]. While these methods capture richer context, they extract each modality feature in isolation and then merge features post hoc, limiting their ability to jointly reason across modalities. Furthermore, lack of unified data space, and the representational gap between e.g. symbolic (textual) and continuous (geospatial) data introduces alignment challenges and interpretability issues in downstream clustering or topic interpretation.
2.3 Graph-based Topic Modeling
Several studies explored graph-based approaches to better capture semantic relationships and improve topic coherence. For example, Wang et al. [wang2015community] constructed word co-occurrence graphs from corpora and applied community detection algorithms, such as the Louvain method, to identify topical clusters, demonstrating superior performance on short text datasets. Similarly, Zhao et al. [zhao2018graph] proposed GSTM, a graph-based short text topic model that utilizes word co-occurrence graphs and community detection to improve topic coherence on sparse data. Beyond purely graph-based clustering, Graph Topic Models (GTM) have integrated Graph Neural Networks (GNNs) with topic modeling to jointly learn from document-word heterogeneous graphs, capturing both network structure and semantic information [zhou2020]. However, despite their ability to incorporate relational structure, GTMs still struggles to fully capture deeper or long-range semantic dependencies, limiting their effectiveness in modeling more abstract or nuanced topics. Earlier work by Newman and Girvan [newman2006finding] established a theoretical connection between community detection in bipartite document-word networks and topic model inference, highlighting the natural synergy between network science and probabilistic topic discovery. Extensions of LDA incorporating graph structures (such as Graph-LDA) have also been proposed to embed word similarity relations directly within the probabilistic framework, yielding more semantically coherent topics [doshi2015]. More recently, Graph2Topic [graph2topic2023] presented a topic modeling framework that leverags graph clustering methods on embedding-based sentence representations. This approach improves the quality and interpretability of topics without requiring a predefined number of topics, demonstrating the practical benefits of graph-embedding integration in topic discovery. Overall, graph-based topic modeling represents a promising direction that exploits the structural properties of the input data, allowing to integrate advances in network science and deep learning to overcome limitations of traditional models.
3 Methodology
Our approach leverages a set of geotagged social media posts, each represented by a semantic embedding $h_{i}$ (derived from the textual content) and a geographic coordinate $l_{i}$. To obtain rich semantic representations, we use Sentence-BERT (SBERT), which fine-tunes transformer-based language models in a Siamese architecture. This structure enables the model to directly optimize for sentence-level similarity, and has been shown in follow-up studies to effectively capture semantic proximity in short, informal texts such as tweets [Reimers2019]. To facilitate joint reasoning over these heterogeneous data types, we convert all modalities into a unified graph structure, serving as a common representation space. Graphs offer a flexible and expressive framework for capturing both local and global relationships within and across modalities. This design enables the neural model to propagate information through neighborhoods in a context-aware and modality-aware manner, something that sequential or flat representations cannot achieve effectively. While this study focuses on semantic and geographic signals, our methodology is not limited to these two modalities; without loss of generality, it can be extended to incorporate additional modalities such as temporal, sentiment, or even visual information, provided they can be represented in a graph-compatible format. We explore two complementary graph construction strategies. In the MonoGraph formulation, we define a unified graph $\mathcal{G}_{u}=(\mathcal{V},\mathcal{E}_{u})$, where edge weights integrate both modalities:
$w_{ij}=\mu\cdot\cos(\mathbf{h}_{i},\mathbf{h}_{j})+\theta\cdot\frac{1}{1+d(\mathbf{l}_{i},\mathbf{l}_{j})},$
where $h_{n}$ are the semantic embeddings and $l_{n}$ are the geographic coordinates, and $\mu$ and $\theta$ control the relative contributions of semantic and geographic similarities, respectively. This fused graph is then processed by a single Graph Convolutional Neural Network (GCN) encoder that learns joint representations capturing both semantic and spatial proximity. In the MultiGraph formulation, we construct two independent graphs over the same node set $\mathcal{V}$: a semantic graph $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E}_{s})$, with edge weights:
$w_{ij}^{(s)}=\cos(\mathbf{h}_{i},\mathbf{h}_{j}),$
and a geographic graph $\mathcal{G}_{g}=(\mathcal{V},\mathcal{E}_{g})$, with edge weights:
$w_{ij}^{(g)}=\frac{1}{1+d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})},$
where $\mathbf{p}_{i}=(\phi_{i},\lambda_{i})$ denotes the latitude–longitude coordinates of node $i$ in degrees under EPSG:4326, and $d_{\mathrm{hav}}$ is the great-circle (Haversine) distance given by:
$d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})=2R\arcsin\left(\sqrt{\sin^{2}\left(\frac{\phi_{j}-\phi_{i}}{2}\right)+\cos(\phi_{i})\cos(\phi_{j})\sin^{2}\left(\frac{\lambda_{j}-\lambda_{i}}{2}\right)}\right),$
where $R$ is the Earth’s radius (e.g., $R\approx 6371\,\mathrm{km}$), and all angular quantities $\phi$ and $\lambda$ are expressed in radians before computation. Similarly, this graph is processed with a dedicated GCN network as well.
Figure 1: Layout of the two models: MonoGraph (a), and MultiGraph (b)
3.1 Network Architectures
A schematic overview of both architectures is provided in Fig. 1. We employ Graph Convolutional Networks (GCNs) to model local context and neighborhood interactions across semantic and geographic dimensions in both architectures. GCNs operate via a localized message-passing scheme, updating node features according to:
$\mathbf{X}^{(l+1)}=\sigma\left(\hat{A}\mathbf{X}^{(l)}\mathbf{W}^{(l)}\right),$
where $\hat{A}$ is the normalized adjacency matrix, $\mathbf{X}^{(l)}$ denotes the feature matrix at layer $l$, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\sigma$ is a non-linear activation function such as ReLU. Our implementation uses two GCN layers with ReLU activations. In the MultiGraph setup, dropout and layer normalization are applied between layers to enhance generalization and training stability. The MonoGraph architecture processes the unified graph with a shared GCN encoder to learn joint node embeddings. Conversely, the MultiGraph architecture builds two separate GCN branches, one for semantic and one for geographic graphs, producing semantic embeddings $\mathbf{x}_{i}^{(s)}$ and geographic embeddings $\mathbf{x}_{i}^{(g)}$ for each node $i$. These embeddings are then fused through a learnable multi-head attention mechanism. In this mechanism, the geographic embeddings serve as the queries $(Q)$, while the semantic embeddings act as both the keys $(K)$ and values $(V)$, enabling the model to selectively attend to relevant semantic features conditioned on geographic context:
$\mathbf{Z}_{i}=\text{Multi-Head}\big(\mathbf{Q}=\mathbf{x}_{i}^{(g)},\mathbf{K}=\mathbf{x}_{i}^{(s)},\mathbf{V}=\mathbf{x}_{i}^{(s)}\big),\quad\mathbf{z}_{i}=\frac{1}{h}\sum_{j=1}^{h}\mathbf{Z}_{i}^{(j)},$
where the multi-head attention layer first linearly projects $(Q,K,V)$ into $h$ subspaces (attention heads) of dimension $d_{k}=\frac{\text{embed\_dim}}{h}$, then computes scaled dot-product attention for each head as
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
and finally concatenates the outputs from all attention heads into a single vector, which is then linearly transformed to produce the fused embedding. Here, $h$ denotes the number of attention heads. This fusion mechanism enables adaptive balancing of semantic and spatial information, selectively modulating the influence of nodes and edges to optimize the overall training objective.
3.2 Loss Design and Training Objective
To guide the learning of node representations, we use a composite loss that combines contrastive, coherence, and alignment objectives. The contrastive loss helps preserve local semantic structure. The coherence loss promotes cluster structure. The alignment loss further stabilizes training by encouraging each node to stay close to the centroid of its assigned cluster. By jointly optimizing these losses, the model is able to learn embeddings that capture both local neighborhood structure and global cluster organization, while also maintaining alignment with cluster prototypes. Clustering is performed using the K-means algorithm with a fixed random seed and fixed cluster numbers to ensure consistency across training epochs, although any other clustering algorithm, differentiable or not, can be used without modifying the core framework.
Coherence Loss: The coherence loss consists of two components: intra-cluster similarity, which ensures topics are internally consistent, and inter-cluster dissimilarity, which discourages overlap between topics.
Intra-cluster Similarity: The intra-cluster similarity measures how similar the embeddings of nodes within the same cluster are to each other. Intuitively, nodes that belong to the same community should have similar representations, as they share common structural or semantic properties. This similarity is commonly quantified using normalized pairwise cosine similarity, which has been shown to be an effective measure for assessing the compactness of clusters, particularly for text and high-dimensional data [ghosh2006]. The cosine similarity captures the angular distance between embeddings, providing a metric that is robust to differences in magnitude while emphasizing the relative similarity of node features:
$\mu_{\text{intra}}=\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)=C(j)\}}}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}$
where $K$ is the number of clusters, and $C_{k}$ denotes the set of nodes in cluster $k$. The denominator normalizes the similarity by the number of node pairs in each cluster. Cosine similarity is computed as:
$\cos(\mathbf{z}_{i},\mathbf{z}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|_{2}\|\mathbf{z}_{j}\|_{2}}$
Here, $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ represent the feature embeddings of nodes $i$ and $j$, respectively, and are normalized by their L2 norms to ensure that only the direction (rather than magnitude) of the vectors influences the similarity measure.
Maximizing intra-cluster similarity encourages the model to learn node embeddings that are close to one another within the same cluster. This promotes compactness and internal consistency in the embedding space, improving the semantic or spatial cohesion of discovered clusters. In this context, higher values of intra-cluster similarity, typically close to 1, indicate tight and cohesive clusters, while values near 0 or -1 suggest poorly defined clusters.
Inter-cluster Similarity: The inter-cluster similarity captures how similar node embeddings are across different clusters. Ideally, embeddings from separate clusters should be dissimilar. We define this as:
$\mu_{\text{inter}}=\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)\neq C(j)\}}}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}$
Here, $P_{\text{inter}}$ denotes all node pairs from distinct clusters. High inter-cluster similarity values would indicate insufficient separation between clusters, while values close to -1 reflect distinct and well-separated clusters. So, ideally we aim for low inter-cluster similarity, with values closer to -1 indicating optimal separation between clusters.
To jointly optimize intra-cluster compactness and inter-cluster separation, we define the coherence loss:
$\mathcal{L}_{\text{coherence}}=-\left(\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}-\lambda\cdot\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}\right)$
where $\lambda$ is a hyperparameter that controls the trade-off between maximizing intra-cluster similarity and minimizing inter-cluster similarity. This loss function combines both clustering principles into a single differentiable objective. The negative sign transforms the coherence score into a minimization problem, consistent with standard optimization procedures. Minimizing $\mathcal{L}_{\text{coherence}}$ leads to embedding spaces where nodes within the same cluster are tightly grouped, while nodes from different clusters are well-separated, enhancing both interpretability and performance of downstream clustering tasks.
Contrastive Loss: To preserve local semantic or structural neighborhoods, we use a contrastive loss that pulls embeddings of similar nodes closer and pushes dissimilar ones apart. For a batch of positive and negative node pairs, the contrastive objective is defined as:
$\mathcal{L}_{\text{contrast}}=-\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{(i,k)\in\mathcal{N}_{i}}\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
where $\mathcal{P}$ denotes the set of positive node pairs (e.g., semantically or spatially close), $\mathcal{N}_{i}$ includes both positives and negatives for node $i$, and $\tau$ is a temperature hyperparameter controlling the sharpness of the distribution. Contrastive loss enforces local consistency in the learned embeddings, ensuring that nodes with similar contexts (semantic or spatial) are embedded close together. This regularizes the representation space and improves generalization, especially in heterogeneous graphs.
Alignment Loss: The alignment loss encourages node embeddings to remain close to the centroid of their assigned cluster. This constraint stabilizes training and reduces noise by enforcing tighter adherence to cluster prototypes. The loss is defined as:
$\mathcal{L}_{\text{align}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{c}_{C(i)}\right\|_{2}^{2}$
where $\mathbf{z}_{i}$ is the embedding of node $i$, and $\mathbf{c}_{C(i)}$ is the centroid of cluster $C(i)$, computed as the mean embedding of all nodes in that cluster. This loss helps to maintain cluster cohesion by anchoring each node close to its corresponding cluster center. When combined with coherence and contrastive losses, alignment loss enhances the structural integrity of the latent space, especially in high-noise environments or early training stages.
Total Loss Objective: To effectively learn cluster-aware and structure-preserving embeddings, we combine all three objectives into a unified loss function. The total loss is defined as:
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{contrast}}+\beta\cdot\mathcal{L}_{\text{coherence}}+\gamma\cdot\mathcal{L}_{\text{align}}$
where $\alpha$, $\beta$, and $\gamma$ are weighting coefficients that control the relative contribution of each component. This composite formulation allows the model to jointly optimize for local consistency, global cluster structure, and prototype alignment.
3.3 Embedded Features Clustering
After obtaining the embedded features produced by the final layer of the GCN encoder, we apply unsupervised clustering to the resulting embeddings $\mathbf{z}_{i}$. Clustering helps to reveal latent topic or community structures in the joint semantic-geographic representation. We use Spectral Clustering, which first computes a similarity matrix that reflects the pairwise relationships between data points, and then utilizes the eigenvectors of the graph Laplacian to partition the data. This method is particularly effective for high-dimensional and non-convex data distributions, such as those found in graph-based embeddings [von2007]. Likewise, we fix the number of clusters $K=10$ to stabilize the process and reduce variability, ensuring consistent topic identification across runs. This choice also enhances reproducibility by maintaining the same clustering configuration.
4 Validation
A comprehensive validation was conducted to rigorously assess the performance and generalization capability of the proposed methodology. We employed a multi-faceted experimental procedure for this purpose. First, we evaluated the models’ ability to capture coherent patterns and embed the social media data into clusters that are cohesive and distinct, corresponding to high intra-cluster similarity and low inter-cluster similarity. Next, we benchmarked their performance against the most comparable method identified in the literature, as well as a sequential baseline, across four diverse datasets. Finally, the utility of the proposed models was demonstrated through a case study on the 2017 Hurricane Harvey in the United States, in which we evaluated the interpretability and the spatial autocorrelation patterns found in the relevant topics, which are proven to be significant for disaster management applications.
4.1 Experimental Setup
Datasets: As the models were designed specifically for the analysis of social media posts in disaster situations, we validated our proposed models using four tweet datasets related to distinct real-world disaster scenarios: (1) the 2014 Napa earthquake [brocher2015], which occurred on August 24, 2014. The geographical bounds focus on the Napa Valley region in California, covering 4,287 km². (2) Hurricane Harvey in 2017 [noaa2017], a Category 4 hurricane that impacted the southeastern Texas region from August 25 to September 7, 2017. The dataset reflects a large spatial extent of 26,736 km², particularly around Houston. (3) The 2021 Ahr Valley flood [koks2021], caused by heavy rainfall, which took place in July 2021. The dataset covers the 1,256 km² area of the Ahr Valley in Germany. (4) The 2023 Chile wildfires [cordero2024], which affected central and southern Chile and were most intense in February 2023. The dataset spans a vast area of 546,905 km². All the datasets were collected using Twitter’s former v1.1 streaming and recent search API endpoints, while filtering by the bounding box Polygon (Geographic Bounds), and the posting time (Time Frame), as depicted in Table 1.
As a common practice in natural language processing (NLP), we preprocessed the raw text data (tweets) before performing further analysis. This preprocessing is done to ensure consistency, remove noise, and improve the quality of the data. Specifically, we removed URLs using a regular expression and converted the text to lowercase for keyword extraction. HTML entities were unescaped, Mentions (i.e., @references) were removed, and extra whitespace was stripped. Non-alphanumeric characters were removed, followed by the elimination of numbers. For BERT-specific cleaning, Mentions were replaced with @user, URLs were converted to http, and newline characters were removed. These steps ensure that the tweets are appropriately formatted for further analysis and model training.
Table 1: Properties of the datasets used for validation. The time frame and geographic bounds were used for filtering. In addition, we computed the number of posts within each dataset and the size of the cover area.
Dataset Posts Time Frame Geographic Bounds Cover Area
Ahr Valley floods 11,177 2021-07-01 – 2021-07-31 (50.01, 6.00, 50.75, 8.00) 1,256 km²
Hurricane Harvey 28,933 2017-08-25 – 2017-09-07 (27.58, –98.77, 30.66, –90.86) 26,736 km²
Chile wildfires 100,707 2023-02-01 – 2023-02-28 (–55.95, –76.89, –17.34, –65.22) 546,905 km²
Napa earthquake 90,907 2014-08-24 – 2014-08-25 (37.11, –123.05, 38.98, –121.04) 4,287 km²
Network Training and Parametrization: The models was trained in an unsupervised manner, without access to ground truth topic or cluster labels, using the compound loss, as detailed in the Section 3.2. Input features consisted of SBERT-generated semantic embeddings for textual content and corresponding geolocation coordinates for graph construction. These modalities were used to define semantic and geographic edges in both the MonoGraph and MultiGraph formulations. Training was performed using the Adam optimizer with an initial learning rate of 1e-3. Models were trained for 50 epochs, with early stopping based on the stability of the coherence score over a held-out validation set. In the MultiGraph setup, we applied a dropout rate of 0.3 and used two Torch GCN layers per branch. The multi-head attention mechanism used for embedding fusion consisted of 4 heads and incorporated layer normalization between fusion stages. To efficiently manage memory, avoid over-sized graphs, and enable consistent training across varying dataset sizes, we adopted a strided chunking protocol. In specific, each dataset was partitioned into multiple non-overlapping chunks using a fixed stride value. Given a stride $s$, each chunk consisted of every $s$-th entry from the original dataset, starting from a different offset. Formally, we generated $s$ subsets $\{\text{chunk}_{i}\}_{i=1}^{s}$ such that:
$\text{chunk}_{i}=\text{Input}_{\text{gdf}}[i::s]$
This approach preserved the overall diversity and temporal structure of the data while enabling the model to process each chunk independently. The protocol ensures that all data points are covered, and total recovery of the original dataset was verified before training. Each chunk was used sequentially as input to the model in a looped training cycle. Hyperparameter tuning was carried out via a random search over a validation subset of the datasets. Ultimately, we set the contrastive, coherence, and alignment loss weights to $\alpha$ = 0.8, $\beta$ = 0.2, and $\gamma$ = 0.1, respectively. These choices are based on a combination of domain knowledge, empirical tuning, and the relative importance of each loss component for the task at hand. For the contrastive loss, we set the temperature parameter $\tau$ = 0.5. All experiments were executed on an NVIDIA H100 GPU.
Table 2: Clustering evaluation results across different datasets using the two models against the input referenced data
Dataset Ahr Valley Floods Hurricane Harvey Chile Wildfires Napa Earthquake
Model Inter Intra Inter Intra Inter Intra Inter Intra
Input 0.432 0.779 0.401 0.693 0.355 0.525 0.319 0.447
MonoGraph 0.046 0.846 0.045 0.817 -0.071 0.777 -0.054 0.774
MultiGraph 0.248 0.853 0.164 0.824 -0.002 0.790 -0.042 0.791
Figure 2: Similarity matrices and their corresponding 3D clustering visualizations (after dimensionality reduction) for a subset of the Napa dataset. The first row (a, b) shows the input embeddings, while the second row (c, d) and third row (e, f) display the results for MonoGraph and MultiGraph models, respectively.
4.2 Clustering Capability Evaluation
To evaluate the models’ capacity for generating semantically coherent representations, we conducted clustering assessments on both the original SBERT input embeddings and the output embedded features produced by each model. Following the methodology outlined in Section 3.2, we computed two key metrics: intra-cluster similarity, reflecting semantic cohesion within clusters, and inter-cluster similarity, indicating the degree of separation between them. Ideally, effective clustering is characterized by high intra-cluster cohesion and low inter-cluster similarity. Table 2 summarizes the results across the four disaster-specific datasets. The input embeddings serve as a baseline and consistently exhibit moderate cohesion but relatively poor separation, suggesting that the raw SBERT representations lack strong semantic structure. Both MonoGraph and MultiGraph yield substantial improvements. Across all datasets, inter-cluster similarity drops significantly, while intra-cluster cohesion improves. MultiGraph achieves the highest cohesion values on all datasets and attains the lowest (i.e., most negative) inter-cluster similarity scores for the Chile wildfires and Napa earthquake datasets, suggesting well-isolated and internally consistent clusters. On the Ahr Valley floods and Hurricane Harvey datasets, MonoGraph shows better (lower) inter-cluster similarity values, although its intra-cluster scores are slightly lower than MultiGraph’s, yet still markedly better than the baseline. These trends are further reinforced by the visualizations in Fig. 2. The input similarity matrix (top-left) exhibit less well-defined structure, whereas the output matrices (middle-left and bottom-left) display sharper contrast and more structured groupings. Likewise, the 3D cluster plots (right) reveal a transition from scattered distributions in the input space to more compact and semantically coherent clusters after model transformation. Low inter-cluster similarity values indicate strong dissimilarity between clusters, reflecting effective cluster separation. In summary, both models, especially MultiGraph, demonstrate a strong ability to reorganize the input features space in ways that enhance clustering quality. This improvement is crucial for downstream tasks such as topic modeling, trend detection, and spatial-semantic analysis.
4.3 Benchmark and Baseline Comparison
The objective of this evaluation was to quantitatively assess the models’ performance and their capability to generate high-quality clusters across diverse disaster scenarios. For benchmarking purposes, we compared our proposed models against the multimodal pipeline presented by Hanny et al. [hanny2025], as well as the baseline sequential workflow introduced in this study. To rule out the effect of time and sentiment features in the multimodal pipeline, we set the sentiment and time weights in the pipeline to zero during the evaluation. To ensure fair comparison, we used the same exact datasets as used in [hanny2025]. For embedding generation, we utilized the English-only all-MiniLM-L12-v2 SBERT model for the 2014 Napa earthquake and 2017 Hurricane Harvey datasets. Conversely, for the 2021 Ahr Valley flood and 2023 Chile wildfires datasets, we adopted the multilingual distiluse-base-multilingual-cased-v1 model, consistent with the approach in [hanny2025]. To ensure a congruent comparison across models, we chose Topic Quality (TQ), a widely used metric in the community. TQ effectively evaluates the coherence and relevance of topics, making it a reliable measure for assessing the quality of topic models. It is computed based on the top 10 keywords per topic and defined as the product of Topic Coherence (TC) and Topic Diversity (TD). We employed a TF-IDF-based keyword extraction method: For each cluster, the corresponding subset of documents was aggregated, and TF-IDF scores were calculated over this localized corpus. The top 10 highest-scoring terms were selected as representative keywords.
To compute the TQ score, we first constructed a document-term matrix using CountVectorizer, then estimated word and word-pair frequencies across the full corpus. For each topic, pairwise Positive Pointwise Mutual Information (PPMI) between all keyword pairs was calculated and normalized by the negative log joint probability to obtain the normalized variant (NPPMI) [bouma2009] [jurafsky2023]. Formally, the normalized PPMI between two words $w_{i}$ and $w_{j}$ is given by
$\text{NPPMI}(w_{i},w_{j})=\frac{\max(0,\log\frac{P(w_{i},w_{j})}{P(w_{i})P(w_{j})})}{-\log P(w_{i},w_{j})},$
where probabilities are estimated from co-occurrence counts over the corpus. The coherence of a topic is the average NPPMI over all $\binom{10}{2}$ word pairs in the topic, and the overall Topic Coherence is the mean across all valid topics. Topic Diversity quantifies how distinct the discovered topics are by measuring the proportion of unique keywords across all topics. It is calculated as
$\text{TD}=\frac{U}{k\times|T|},$
where $U$ denotes the number of unique words in the top-$k$ keywords across all topics, $k$ is the number of keywords per topic, and $|T|$ is the total number of topics. This penalizes repeated keywords across topics and rewards diversity in word usage. Finally, Topic Quality is computed as the product of the two metrics:
$\text{TQ}=\text{TC}\times\text{TD},$
which favors models that produce both semantically coherent and lexically diverse clusters. All metrics are computed using the same preprocessing pipeline for consistency across evaluated methods.
4.4 TQ Results and Analysis
The Topic Quality (TQ) scores reported in Table 3 clearly demonstrate the effectiveness of our proposed models in generating high-quality topic clusters across diverse disaster-related datasets. The MultiGraph consistently achieves the highest TQ scores in three out of four scenarios, the Ahr Valley floods (0.415), Hurricane Harvey (0.384), and the Napa earthquake (0.195), demonstrating its robustness and adaptability to datasets with varying linguistic, geographic, and topical characteristics. This superior performance can be attributed to its cross-attention mechanism, which enables selective regulation of individual modalities and more effective capture of both global semantic patterns and local topical signals. On the Chile wildfires dataset, however, the MonoGraph slightly outperforms the MultiGraph (0.169 vs. 0.161), with both models performing comparably well overall. This suggests that while the MultiGraph generally excels, the benefits of increased model complexity may vary depending on the nature and quality of the data. In particular, contexts characterized by fewer high-quality labeled signals or more multilingual noise may favor the simpler and more direct MonoGraph architecture, which consistently outperforms baseline methods across all datasets and yields stable, focused clusters despite its relative simplicity.
Table 3: Topic quality scores (TQ) for outputs from our two models (MonoGraph and MultiGraph), the method by Hanny et al. [hanny2025], and the baseline sequential workflow across four datasets.
Dataset MonoGraph MultiGraph Hanny et al. [hanny2025] Sequential Workfollow
Ahr Valley Floods 0.383 0.415 0.235 0.029
Hurricane Harvey 0.368 0.384 0.222 0.042
Chile Wildfires 0.169 0.161 0.117 0.034
Napa Earthquake 0.191 0.195 0.143 0.030
The multimodal pipeline [hanny2025] shows limited performance (TQ ranging from 0.117 to 0.235), likely due to the removal of sentiment and temporal features during evaluation. These components, while integral to the original design, appear essential to the pipeline’s effectiveness. While the model presented in this paper was originally designed to incorporate sentiment and temporal features, these were deliberately neutralized in our evaluation to isolate textual topic modeling performance. The resulting drop in quality highlights the fact that the pipeline’s heavy reliance on auxiliary modalities may limit its effectiveness in settings where such features are unavailable, noisy, or intentionally excluded. The sequential workflow performs the worst overall, with TQ scores falling below 0.05 across all datasets. This approach suffers from treating the stages of the modeling pipeline, embedding, clustering, and keyword extraction, as independent components, which results in fragmented topic structure and poor alignment between cluster semantics and extracted keywords. The absence of feedback or interaction between stages appears to hinder its ability to produce globally meaningful topic representations.
Figure 3: Topic-word distributions for a subset of the Hurricane Harvey dataset, applying MonoGraph (top) and the MultiGraph (bottom) models.
5 Case Study Evaluation
A key motivation behind the design of our models was to ensure interpretability for real-world applications. To complement the experiments in Section 4, we conducted a case study to showcase their utility in a disaster management context. We used a subset of Hurricane Harvey dataset, a tropical storm that occurred in 2017 and hit the state of Texas in the USA on August 25 [noaa2017]. The dataset comprises 28,933 geo-located tweets posted within and around the area of Texas, obtained by filtering tweets based on Geographic Bounds defined by the polygon $\text{Polygon}(-98.77,27.58;-90.86,27.58;-90.86,30.66;-98.77,30.66;-98.77,27.58)$ and the Time Frame from August 25 to September 7, 2017.
5.1 Topic Interpretation
To gain qualitative insights into the latent structure of the discovered clusters, we analyzed both the semantic coherence and spatial distribution of topics generated by the MonoGraph and MultiGraph models applied to a subset of the dataset. The topic-word distributions, summarized in Fig. 3, reveal semantically coherent and lexically diverse clusters that reflect key aspects of disaster-related discourse, and several topics capture geographically grounded and event-focused narratives. In particular, Topic 0 groups terms such as “houston,” “texas,” “harvey,” “hurricaneharvey,” “hurricane” and “flood,” reflecting a strong spatial anchoring of disaster-related conversations. As reflected in the coresponding spatial visualization in Fig. 4, the distribution of this topic closely follows the hurricane’s path, with dense tweet concentrations in Houston and along the Texas coastline. This alignment suggests that the models successfully identify localized, event-driven public responses directly associated with Hurricane Harvey’s impact zone. Another distinct cluster is Topic 2 in MonoGraph and Topic 1 and in MultiGraph, centered on traffic-related concerns, including terms like “traffic,” “accident,” “mins,” “delay,” and “stop.” the spatial maps reveal concentrated tweet activity along major transport corridors in southeastern Texas, likely corresponding to road closures, congestion, and transportation disruptions resulting from flooding and evacuation efforts.
Figure 4: Spatial distribution (KDE) of the Topic-words for a subset of the Hurricane Harvey dataset, applying MonoGraph (top) and the MultiGraph (bottom) models.
Other topics represent broader or more indirect engagement with the event. Topic 8 in MonoGraph and Topic 5 and in MultiGraph, captured meteorological observations, with keywords such as “temperature,” “humidity,” “clouds,” “rain,” and “wind,” are more evenly distributed across central Texas, including Austin and San Antonio. These spatial patterns suggest contributions from users outside the most severely affected areas, often in the form of commentary or weather updates rather than first-hand impact reports. Weather-related discourse is further reflected in Topics 2, 7, 5 and 9, in MonoGraph and Topics 3, 4, 9, 6 and 8 in MultiGraph, featuring overlapping terms such as “pressure,” “gust,” “humidity,” “wind,” and “clouds.” While this lexical overlap reduces topic distinctiveness, it underscores the recurring presence of weather-related signals across the corpus, a common characteristic in disaster-related social media data. This also illustrates the trade-off of fixing the number of topics ($K=10$), which facilitates stability but can lead to partial redundancy. Finally, content-sharing and localized observational behavior emerge as a consistent pattern. For example, certain clusters contain terms such as “posted,” “photo,” and place names like “meyerland,” “katy,” and “downtown,” suggesting the dissemination of real-time visual documentation and location-tagged updates by affected residents. This pattern reflects well-known crisis communication behavior on social media, where users share on-the-ground images to inform others and raise awareness.
Figure 5: Moran’s I plots for MonoGraph (top) and MultiGraph (bottom) Models.
5.2 Global Autocorrelation Interpretation
To explore the global distribution of tweet activity associated with different topics, we applied Moran’s I analysis. This method is particularly beneficial as it helps to identify spatial autocorrelation, revealing whether tweet activity related to specific topics is clustered in certain geographic regions or distributed randomly around. The metric ranges from -1 (perfect dispersion) to +1 (perfect clustering), with 0 indicating randomness. A p-value << 0.05 suggests the observed spatial pattern is statistically significant and unlikely due to chance. The related scatter plots and their associated statistics, demonstrated in Fig. 5 complement the spatial intensity patterns visualized in Fig. 4. In the MultiGraph model, Topic 0 shows a Moran’s I of 0.076 (p = 0.001), suggesting weak but statistically significant clustering. Tweets on this topic tend to occur more frequently near other tweets discussing the same theme than by chance, although the clustering is not very strong. The slight clustering might indicate localized interest or discussions related to this topic in specific regions. Topic 2 has a Moran’s I of 0.178 (p = 0.001), indicating weak but significant spatial clustering. The positive value suggests a tendency for tweets on this topic to be more closely located in certain regions. This could reflect a localized trend where users in particular areas are more engaged with or affected by the theme of the topic. Topic 5 has a Moran’s I of -0.032 (p = 0.001), a slightly negative Moran’s I indicating weak spatial dispersion, with tweets being more likely to occur near unrelated content. This suggests a dispersed pattern of topic-related tweets, though the effect is weak. The scattered distribution might indicate that this topic is more widespread, with no strong geographic concentration.
In the MonoGraph model, Topic 0 shows a Moran’s I of 0.065 (p = 0.001), which is similar to the MultiGraph model, showing weak clustering, though the clustering effect is slightly weaker than in the MultiGraph, with tweets on this topic being somewhat more scattered. This could imply a broader discussion, not concentrated in specific areas, though still more frequent in certain locations compared to random. Topic 1 has a Moran’s I of 0.201 (p = 0.001), indicating moderate clustering. Tweets on this topic show a clear tendency to be geographically concentrated in certain regions, which suggests a more localized theme. The higher Moran’s I value suggests that the topic is more likely to spark concentrated discussions within specific areas, potentially reflecting regional events or discussions. Topic 8 has a Moran’s I of -0.035 (p = 0.001), indicating a weakly dispersed distribution of tweets, showing no significant clustering and pointing to random spatial dispersion. The lack of clustering suggests that the topic has less geographic relevance or is more universally spread, with little concentration in specific regions.
Overall, most topics exhibit weak spatial clustering, as indicated by positive but modest Moran’s I values for topics 0, 1, and 2, reflecting slight regional concentration without distinct spatial boundaries. Topic 0, directly related to disasters, shows some geographic focus, while topics 1 and 2, indirectly related to disaster events, display weaker clustering. In the MonoGraph model, Topic 1 demonstrates stronger clustering (Moran’s I = 0.201), suggesting a more localized concentration. Topics 5 and 8 show weak spatial dispersion, with low or negative Moran’s I values, indicating a more even distribution across geographic regions and a lack of clear spatial patterns.
Figure 6: Spatial autocorrelation analysis (the Gi* maps (top) and LISA maps (bottom)) for the MonoGraph model.
Figure 7: Spatial autocorrelation analysis (the Gi* maps (top) and LISA maps (bottom)) for the MultiGraph model.
5.3 Spatial Autocorrelation Analysis
To further investigate the spatial (local) dynamics of identified topics, we applied the Getis-Ord Gi* statistic and Local Indicators of Spatial Association (LISA) to both the MonoGraph outputs. From a disaster management perspective, combining Gi* with LISA provides a powerful way to interpret social media–derived topics in both their intensity and spatial context. Gi* hotspots highlight areas of concentrated disaster-related activity, helping to quickly pinpoint zones of immediate impact for targeted relief. LISA adds a layer of context by classifying regions into high–high, low–low, high–low, and low–high clusters, showing not only where activity is high but also how it relates to surrounding areas. This combined approach enables planners to identify both core impact zones and peripheral but relevant areas, anticipate spillover effects, and spot isolated needs. In our results, the MultiGraph (Fig. 7) tends to produce more localized clusters, which may be better suited for tactical, on-the-ground coordination, while the MonoGraph model’s broader patterns (Fig. 6) can inform strategic, statewide planning. For Topic 0 in both models, representing core Hurricane Harvey discourse, the Gi* analysis revealed statistically significant hotspots spanning Houston, the Texas Gulf Coast, and inland counties along the storm’s path, with corresponding coldspots in northern and western Texas. LISA clustering confirmed persistent high–high (HH) clusters in these impact zones, indicating strong spatial concentration of disaster-related conversation. Such patterns can guide response agencies toward areas most actively engaged in reporting severe impacts, closely aligning with ground-truth disaster footprints.
As it can be seen in Figs. 6 and 7, traffic-related topics (MonoGraph Topic 2, MultiGraph Topic 1) produced more localized hotspots in and around metropolitan Houston and along major transport corridors. Gi* maps highlight narrowly bounded high-value clusters, while the LISA maps show HH concentrations surrounded by low–low (LL) areas, indicating geographically confined yet intense transportation disruptions. This suggests that mobility restoration efforts could be prioritized for specific corridors rather than across broad regions. Weather observation topics (MonoGraph Topic 8 MultiGraph Topic 5) exhibited scattered hotspots interspersed with wide coldspot zones. LISA classifications revealed fewer HH clusters but a notable presence of low–high (LH) outliers, counties with high weather-related discourse surrounded by low-activity neighbors. These may represent isolated monitoring stations, amateur weather observers, or unaffected communities contributing updates, which can complement official meteorological networks during disaster response.
6 Discussion and Limitations
Across all datasets, both MonoGraph and MultiGraph consistently outperformed baseline methods, achieving higher intra-cluster cohesion and lower inter-cluster similarity. Their relative strengths varied with event characteristics, suggesting sensitivity to spatial scale, discourse patterns, and data quality. Fixing the number of clusters at $K=10$ facilitated direct cross-model comparisons and system stability, but also introduced constraints. In some cases, semantically similar clusters emerged, particularly in weather-related topics, where recurring terms such as “pressure,” “humidity,” “wind” were dispersed across multiple clusters. This effect may be amplified in datasets containing unfiltered automated “weather bot” content. Empty or low-content clusters occasionally appeared due to sparse coverage, noise, or embedding collapse. The spatial distribution of events significantly shaped clustering performance. MultiGraph produced the lowest inter-cluster similarity for geographically concentrated events (Chile wildfires, Napa earthquake), benefiting from modality-specific graph encoding and cross-attention fusion. MonoGraph showed better separation in smaller or noisier AOIs (Ahr Valley floods, Hurricane Harvey), where its unified representation was more robust to diffuse or overlapping discourse. Spatially, MultiGraph generated compact, well-defined hotspots with fewer outliers, aligning clusters closely with localized disaster impacts, an asset for targeted response planning. MonoGraph captured broader spatial patterns, useful for representing large-scale narratives but with a tendency to overestimate geographic extent.
Several limitations merit consideration. First, the framework’s performance is sensitive to the quality and balance of input modalities. Second, the current implementation does not explicitly model temporal dynamics, which constrains the ability to capture and analyze the evolution of topics over time. Third, the use of a fixed number of clusters ($K$) may not reflect the true topical complexity of diverse events, suggesting the need for adaptive or non-parametric clustering strategies. Finally, MultiGraph’s scalability to real-time, high-volume streams is untested. These constraints may hinder applicability to scenarios involving extremely high-frequency data or expansive geographic coverage unless complemented by distributed computing or streaming graph processing frameworks. Despite these constraints, both architectures effectively detect geographically specific and semantically coherent topics that align with real-world hazard footprints, offering practical value for enhancing situational awareness and supporting decision-making in disaster management.
7 Conclusion
This work presented an unsupervised multimodal graph-based methodology for the semantic and spatial analysis of geo-social media data. By jointly modeling semantic embeddings and geographic coordinates through graph-based architectures, the approach addresses limitations of prior models that treat these modalities in isolation. Two architectural variants, MonoGraph and MultiGraph, were introduced to accommodate differing data characteristics and application needs. A composite loss function integrating contrastive, coherence, and alignment terms was key to learning robust, cluster-aware representations that preserve both local and global structure. Extensive evaluation on four real-world disaster datasets demonstrated consistent improvements over state-of-the-art baselines in topic coherence, spatial compactness, and overall topic quality. The MultiGraph variant, employing modality-specific encoding with attention-based fusion, achieved the strongest performance in most settings, while qualitative analysis of Hurricane Harvey (case study) confirmed the interpretability and geographic relevance of the discovered topics. The proposed methodology offers a scalable, principled approach to integrate symbolic and spatial signals for geo-social media analysis. Although the current implementation focuses on text and location, it can be extended to additional modalities such as sentiment, temporal dynamics, or imagery, enabling richer, more context-aware analyses in domains including disaster management, urban planning, and public health surveillance. Future directions include incorporating time-aware modeling, improving scalability for real-time data streams, and leveraging social network structures to enhance community-level insights. By unifying multimodal signals in an end-to-end learning model, this work advances toward more holistic and interpretable models for understanding the complex information landscapes of social media.

### v2

@@
 1 Introduction
-Social media platforms continuously generate vast amounts of data, including posts and comments, likes, shares, images, and videos.
-This content is rich, diverse, and real-time, offering a valuable resource for understanding how people interact, express opinions, form communities, and respond to events.
-The widespread adoption of smartphones with built-in location-tracking capabilities has added another dimension to this landscape: geographic context.
-By enabling users to share their locations, these devices have facilitated the integration of geospatial data into online content, enhancing the contextual depth of social media analysis.
-A growing body of research underscores the importance of geo-location data in social media research.
-Prior studies have demonstrated its utility across a wide range of applications, from disaster response [goodchild2010crowdsourcing] to public health monitoring [Signorini2011].
-In particular, geo-referenced content from micro-blogging platforms has been leveraged to model and monitor phenomena such as earthquakes [Havas2021Portability], floods [Huang2018Flood], refugee movements [Havas2021Refugee], and disease outbreaks [Stolerman2023Disease].
-Previous social media research has often been focusing on semantic topic modeling, a fundamental analytical algorithmic technique to uncover latent thematic structures within large volumes of user-generated text.
-A range of scholarly works have utilized topic modeling to reveal latent thematic structures within such texts [Resch2018] [Hannigan2019].
-However, the resulting topic representations are rarely contextualized with additional data modalities.
-Early efforts to analyze geo-social textual media in a multimodal setting typically employed sequential workflows, wherein each modality was processed in isolation through distinct modules [yin2011geographical].
-While these methods have demonstrated utility in various applications, their modular and linear design often hinders the ability to capture complex interdependencies between modalities.
-Furthermore, this sequential structure can introduce dependencies between processing stages, rendering the overall output sensitive to the order in which analyses are performed.
-More recent approaches introduced advancements toward integrating multiple modalities such as time, geographic space, semantic topics, and sentiments in the analysis of social media [gonzalez2024neural] [hanny2025].
-However, they still depend on numerous isolated components and fragmented intermediate processing steps, falling short of delivering a truly end-to-end processing methodology.
-Nonetheless, inherent heterogeneity of input modalities in this case has already added to the complexity of the task.
-For example, text is symbolic and sequential, while location data is numeric and geospatial, often resulting in misaligned feature representations and inconsistencies during fusion.
-The lack of a unified machine learning model constrains cross-modal interactions and introduces fragile interdependencies, whereby variations in preprocessing or processing order can significantly impact both performance and interpretability.
-This paper presents a new methodology for multimodal geospatial machine learning, that jointly embeds semantic and geospatial information in a unified graph-based space, using end-to-end learning with a composite loss to generate clusters that are both semantically coherent and spatially interpretable.
-The proposed methodology is application-agnostic and can, without loss of generality, be adapted to diverse forms of multimodal data integration, making it well-suited for a wide range of data fusion scenarios and downstream tasks across domains.
-To capture cross-modal relationships in a holistic and structured manner, we construct a unified multimodal embedding space based on a graph structure, where we can adapt diverse data types, including text, geolocation, and optionally emotions or temporal signals, into a shared representation space.
-In this formulation, nodes in the graph represent user-generated post embeddings, while edges encode relational signals such as semantic similarity, geographic proximity, or user relations.
-This enables the model to attend to multimodal interactions in a context-aware fashion.
-At the core of our methodology, we train an artificial neural network in an unsupervised manner to aggregate information from a node’s neighbors, where a modified message passing mechanism allows scaling by the corresponding edge weights.
-A key innovation of our model is the use of a composite loss function inspired by reference-less clustering evaluation approaches.
-To enable robust and discriminative representation learning, the loss integrates contrastive, coherence, and alignment objectives.
-The contrastive component preserves local semantic structure by pulling similar nodes closer to each other in the embedding space while pushing dissimilar ones apart.
-The coherence objective encourages the formation of well-defined clusters by maximizing intra-cluster similarity and inter-cluster separation.
-At the same time, the alignment term stabilizes training by ensuring each node remains close to the centroid of its assigned cluster, thus enhancing representational consistency.
-By embedding these objectives directly into the learning process, the model is able to discover clusters that are both semantically meaningful and spatially coherent.
-Our approach is specifically tailored to support disaster management efforts.
-In emergencies triggered by natural hazards such as floods, geo-social media provides rich multimodal data from affected communities, enhancing situational awareness for first responders.
-Through a case study of the 2017 Hurricane Harvey in the USA, we demonstrate how geographically fine-grained multimodal clusters of social media posts closely align with the disaster’s observed impacts, illustrating the practical value of our method in real-world crisis scenarios.
-The framework itself is designed to be application-agnostic and, without loss of generality, can be adapted to various forms of multimodal data integration.
-It is well-suited for diverse data fusion scenarios and downstream tasks across different domains.
-Empirical evaluations demonstrate that our approach outperforms existing baselines in topic coherence, spatial compactness, and interpretability, providing a scalable and principled framework for multimodal social media analysis.
-In summary, this work makes the following distinct contributions:
-  • Introducing a multimodal graph-based framework that seamlessly integrates multiple modalities (specifically semantic and geospatial data) into a unified representation space.
-  • Proposing a domain-agnostic methodology for multimodal data analysis, underpinned by two unsupervised architectures, MonoGraph and MultiGraph, that adapt to diverse data characteristics and application needs, positioning the approach as a versatile framework for research and practice in fields such as disaster management, public health, and social media analytics.
-  • Developing a composite loss function that combines contrastive, coherence, and alignment objectives, ensuring the generation of semantically coherent and spatially interpretable clusters.
-  • Demonstrating the effectiveness of our approach through a real-world disaster case study, where geographically fine-grained multimodal clusters of social media posts closely align with observed impacts.
+Social media platforms continuously generate vast volumes of user-generated content, including posts, comments, images, videos, and interaction signals such as likes and shares.
+This content is rich, diverse, and real-time, offering a valuable resource for understanding how people communicate, express opinions, form communities, and respond to events.
+The widespread adoption of smartphones equipped with location-tracking capabilities has added an important additional dimension to this landscape: geographic context.
+By enabling users to share their locations, these devices have facilitated the integration of geospatial information into online content, significantly enhancing the contextual depth of social media analysis.
+A growing body of research has demonstrated the value of geo-referenced social media data across a wide range of applications, including disaster response [goodchild2010crowdsourcing], public health monitoring [Signorini2011], and crisis management.
+In particular, location-aware content from microblogging platforms has been leveraged to study and monitor phenomena such as earthquakes [Havas2021Portability], floods [Huang2018Flood], refugee movements [Havas2021Refugee], and disease outbreaks [Stolerman2023Disease].
+These studies highlight the potential of combining textual and spatial signals to gain timely and actionable insights from social media streams.
+Much of the existing social media analysis literature has focused on semantic topic modeling as a fundamental technique for uncovering latent thematic structures within large collections of user-generated text [Resch2018, Hannigan2019].
+While effective in capturing semantic regularities, these approaches typically treat text in isolation and rarely integrate additional modalities in a principled manner.
+Early efforts toward multimodal geo-social analysis often relied on sequential or modular workflows, in which each modality (e.g., text, location, time) was processed independently before being combined at a later stage [yin2011geographical].
+Such designs limit the ability to capture complex cross-modal interactions and often introduce brittle dependencies between processing stages, making results sensitive to preprocessing choices and execution order.
+More recent approaches have advanced toward integrating multiple modalities, including temporal signals, geographic space, semantic topics, and sentiment [gonzalez2024neural, hanny2025].
+However, these methods still depend on fragmented pipelines with multiple isolated components, falling short of truly end-to-end multimodal learning.
+The inherent heterogeneity of input modalities further complicates this problem: text is symbolic and sequential, while geographic data is continuous and spatial.
+As a result, feature representations across modalities are often misaligned, leading to inconsistencies during fusion and reduced interpretability.
+The absence of a unified learning framework constrains cross-modal reasoning and limits the model’s ability to jointly exploit semantic and spatial structure.
+In this paper, we propose a unified multimodal graph-based learning framework that jointly embeds semantic and geographic information into a shared representation space using end-to-end training.
+Our methodology constructs parallel semantic and geographic graphs over social media posts and learns representations that are both semantically coherent and spatially interpretable.
+Nodes in the graph correspond to post-level embeddings, while edges encode relational signals such as semantic similarity and geographic proximity.
+This formulation enables the model to reason over multimodal interactions in a structured and context-aware manner, rather than relying on post hoc fusion of independently learned features.
+At the core of our approach is an unsupervised graph neural network that aggregates information from each node’s neighborhood through a weighted message-passing mechanism.
+A key innovation of the proposed framework is the use of a composite loss function inspired by reference-less clustering evaluation principles.
+The loss integrates three complementary objectives: a contrastive loss that preserves local semantic neighborhoods by pulling similar nodes closer while pushing dissimilar ones apart; a coherence loss that promotes compact and well-separated clusters; and an alignment loss that stabilizes training by encouraging consistency between node embeddings and their assigned cluster centroids.
+By embedding these objectives directly into the learning process, the model discovers topic clusters that are simultaneously semantically meaningful and geographically grounded.
+Our approach is motivated by applications in disaster management, where geo-social media provides rich, fine-grained information from affected communities during emergencies.
+Through a case study of the 2017 Hurricane Harvey event in the United States, we demonstrate that the proposed framework produces geographically localized and semantically coherent topic clusters that closely align with observed disaster impacts.
+While this case study illustrates the practical utility of the model in crisis scenarios, the framework itself is application-agnostic and can be readily adapted to other multimodal data integration tasks across domains such as public health, urban analytics, and social sensing.
+In summary, this work makes the following contributions:
+  • We introduce a unified multimodal graph-based framework that integrates semantic and geospatial information into a shared representation space for unsupervised topic discovery.
+  • We propose two unsupervised architectures, MonoGraph and MultiGraph, that flexibly support different levels of multimodal interaction while remaining domain-agnostic.
+  • We develop a composite loss function combining contrastive, coherence, and alignment objectives to generate clusters that are both semantically coherent and spatially interpretable.
+  • We validate the effectiveness of the proposed approach through a real-world disaster case study, demonstrating improved topic quality, spatial compactness, and interpretability compared to existing baselines.
 2 Related work
@@
 2.1 Topic Modeling
-Topic modeling is a widely used technique for discovering latent semantic topics within large document collections, serving as a variant of automated content analysis that assigns one or more “topic codes” to each document, thereby providing a structured thematic summary [hoyle2022content].
+Topic modeling is a widely used technique for discovering latent semantic topics within large document collections, providing a structured thematic summary by assigning one or more topic labels to each document [hoyle2022content].
 Among the foundational methods, Latent Dirichlet Allocation (LDA) [blei2003latent] introduced a fully generative probabilistic framework in which each document is modeled as a mixture over latent topics, and each topic is represented as a multinomial distribution over words.
-The generative idea of LDA has been extended to incorporate additional dimensions such as sentiment analysis [Lin2009Joint], temporal dynamics [Blei2006Dynamic], and social media user information [zhao2011comparing], enriching the interpretability and applicability of topic models across domains.
+The LDA framework has been extended to incorporate additional dimensions such as sentiment [Lin2009Joint], temporal dynamics [Blei2006Dynamic], and user-level information in social media [zhao2011comparing], broadening the applicability of topic models across domains.
 However, LDA and its variants fundamentally rely on the bag-of-words assumption, which ignores word order and semantic relationships between words, limiting their effectiveness in handling short, noisy, or highly contextual texts such as tweets and social media posts [hong2010empirical].
@@
 This method utilizes techniques like Uniform Manifold Approximation and Projection (UMAP) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) for dimensionality reduction and flexible clustering.
-Grootendorst [grootendorst2022bertopic] further advanced this line of work with BERTopic, which incorporates transformer-based contextual embeddings (e.g.
-from BERT) combined with refined clustering strategies, consistently outperforming traditional LDA and Top2Vec in both coherence and interpretability.
-Despite their progress, these models generally follow a sequential and fragmented framework and often struggle to balance semantic accuracy, interpretability, and scalability, particularly when applied to the diverse and noisy nature of social media data.
+Grootendorst [grootendorst2022bertopic] further advanced this line of work with BERTopic, which combines transformer-based contextual embeddings with refined clustering strategies, consistently outperforming traditional LDA and Top2Vec in both coherence and interpretability.
+Complementary empirical evidence from related domains further highlights the limitations of purely embedding-driven approaches.
+In particular, studies of pre-trained language models in simple knowledge graph question answering (KGQA) show that, while contextual embeddings capture strong lexical and semantic regularities, they struggle to reliably encode explicit relational structure without dedicated graph-based inductive bias [hu2023empirical].
+These findings suggest that high-quality semantic embeddings alone are often insufficient for tasks requiring structured relational reasoning, motivating approaches that integrate neural language representations with explicit graph modeling.
 2.2 Spatio-semantic Topic Modeling
@@
 While these methods capture richer context, they extract each modality feature in isolation and then merge features post hoc, limiting their ability to jointly reason across modalities.
-Furthermore, lack of unified data space, and the representational gap between e.g.
-symbolic (textual) and continuous (geospatial) data introduces alignment challenges and interpretability issues in downstream clustering or topic interpretation.
+The absence of a unified representation space, together with the gap between symbolic textual data and continuous geospatial representations, further introduces alignment and interpretability challenges in downstream clustering or topic interpretation.
 2.3 Graph-based Topic Modeling
-Several studies explored graph-based approaches to better capture semantic relationships and improve topic coherence.
-For example, Wang et al.
-[wang2015community] constructed word co-occurrence graphs from corpora and applied community detection algorithms, such as the Louvain method, to identify topical clusters, demonstrating superior performance on short text datasets.
+Graph-based topic modeling has emerged as an effective alternative to traditional probabilistic models by explicitly exploiting relational structure to improve semantic coherence, particularly for short and sparse texts.
+Early approaches construct word co-occurrence graphs from text corpora and apply community detection algorithms to identify topical clusters.
+For instance, Wang et al.
+[wang2015community] demonstrated that community detection methods such as Louvain can outperform Latent Dirichlet Allocation (LDA) on short text datasets.
 Similarly, Zhao et al.
-[zhao2018graph] proposed GSTM, a graph-based short text topic model that utilizes word co-occurrence graphs and community detection to improve topic coherence on sparse data.
-Beyond purely graph-based clustering, Graph Topic Models (GTM) have integrated Graph Neural Networks (GNNs) with topic modeling to jointly learn from document-word heterogeneous graphs, capturing both network structure and semantic information [zhou2020].
-However, despite their ability to incorporate relational structure, GTMs still struggles to fully capture deeper or long-range semantic dependencies, limiting their effectiveness in modeling more abstract or nuanced topics.
-Earlier work by Newman and Girvan [newman2006finding] established a theoretical connection between community detection in bipartite document-word networks and topic model inference, highlighting the natural synergy between network science and probabilistic topic discovery.
-Extensions of LDA incorporating graph structures (such as Graph-LDA) have also been proposed to embed word similarity relations directly within the probabilistic framework, yielding more semantically coherent topics [doshi2015].
-More recently, Graph2Topic [graph2topic2023] presented a topic modeling framework that leverags graph clustering methods on embedding-based sentence representations.
-This approach improves the quality and interpretability of topics without requiring a predefined number of topics, demonstrating the practical benefits of graph-embedding integration in topic discovery.
-Overall, graph-based topic modeling represents a promising direction that exploits the structural properties of the input data, allowing to integrate advances in network science and deep learning to overcome limitations of traditional models.
+[zhao2018graph] proposed the Graph-based Short Text Topic Model (GSTM), which leverages word co-occurrence graphs to alleviate data sparsity and improve topic coherence.
+Subsequent work has sought to integrate graph structure more deeply into topic modeling.
+Newman and Girvan [newman2006finding] established a theoretical connection between community detection in bipartite document–word networks and topic model inference, highlighting the natural synergy between network science and probabilistic topic discovery.
+Extensions of LDA incorporating graph structure, such as Graph-LDA [doshi2015], embed word similarity relations directly into the generative framework, yielding more semantically coherent topics.
+More recently, Graph Topic Models (GTMs) have combined Graph Neural Networks (GNNs) with document–word heterogeneous graphs to jointly learn semantic representations and relational structure [zhou2020].
+While these models benefit from relational inductive biases, they often struggle to capture deeper or long-range semantic dependencies and remain limited in modeling more abstract or nuanced topics.
+Embedding-based graph clustering approaches further extend graph-based topic modeling by operating directly on learned sentence or document representations.
+Graph2Topic [graph2topic2023], for example, applies graph clustering to embedding-based representations, improving topic interpretability and quality without requiring a predefined number of topics.
+This line of work demonstrates the practical advantages of combining graph structure with embedding-based semantic representations for topic discovery.
+In parallel, recent advances in generic graph representation learning have introduced self-supervised contrastive objectives for learning node or graph embeddings without labels.
+GraphCL [you2020graphcl] performs graph-level contrastive learning by contrasting different augmented views of the same graph, while MVGRL [hassani2020] learns node representations by contrasting multiple structural views of a graph.
+Although these methods achieve strong performance on generic graph learning benchmarks, they are not designed for topic discovery and do not explicitly model semantic clustering, multimodal fusion, or geographically grounded objectives.
+Beyond topic modeling, heterogeneous network embedding methods have been widely studied in domains such as recommendation systems, where multiple node and edge types are jointly modeled within a unified graph representation [shi2016heterogeneous, zhang2018recommendation].
+These methods learn relational embeddings from heterogeneous graphs but are typically optimized for supervised or semi-supervised prediction objectives, rather than unsupervised topic discovery.
+As a result, they do not explicitly optimize for cluster-level semantic coherence, spatial grounding, or interpretability, which are central to geo-social topic modeling.
+Across topic modeling, spatio-semantic extensions, and graph-based representation learning, prior work has largely addressed semantic structure, spatial information, and relational modeling in isolation.
+Classical and embedding-based topic models focus on semantic similarity, while spatially grounded topic models typically introduce geographic information as auxiliary metadata or enforce it as a post hoc constraint on semantically derived clusters [Yuan2013Who, Hong2012Discovering].
+In parallel, recent graph-based multimodal fusion and contrastive learning methods emphasize learning robust relational embeddings, but are generally not designed for clustering-oriented topic discovery.
+Consequently, semantic structure, geographic proximity, and clustering objectives are rarely optimized jointly within a single end-to-end framework.
+Our framework addresses this gap by explicitly modeling semantic and geographic relations as parallel graph structures and integrating them within a unified learning objective tailored for unsupervised topic discovery, yielding clusters that are both semantically coherent and geographically grounded.
 3 Methodology
-Our approach leverages a set of geotagged social media posts, each represented by a semantic embedding $h_{i}$ (derived from the textual content) and a geographic coordinate $l_{i}$.
-To obtain rich semantic representations, we use Sentence-BERT (SBERT), which fine-tunes transformer-based language models in a Siamese architecture.
-This structure enables the model to directly optimize for sentence-level similarity, and has been shown in follow-up studies to effectively capture semantic proximity in short, informal texts such as tweets [Reimers2019].
-To facilitate joint reasoning over these heterogeneous data types, we convert all modalities into a unified graph structure, serving as a common representation space.
-Graphs offer a flexible and expressive framework for capturing both local and global relationships within and across modalities.
-This design enables the neural model to propagate information through neighborhoods in a context-aware and modality-aware manner, something that sequential or flat representations cannot achieve effectively.
-While this study focuses on semantic and geographic signals, our methodology is not limited to these two modalities; without loss of generality, it can be extended to incorporate additional modalities such as temporal, sentiment, or even visual information, provided they can be represented in a graph-compatible format.
-We explore two complementary graph construction strategies.
-In the MonoGraph formulation, we define a unified graph $\mathcal{G}_{u}=(\mathcal{V},\mathcal{E}_{u})$, where edge weights integrate both modalities:
-$w_{ij}=\mu\cdot\cos(\mathbf{h}_{i},\mathbf{h}_{j})+\theta\cdot\frac{1}{1+d(\mathbf{l}_{i},\mathbf{l}_{j})},$
-where $h_{n}$ are the semantic embeddings and $l_{n}$ are the geographic coordinates, and $\mu$ and $\theta$ control the relative contributions of semantic and geographic similarities, respectively.
-This fused graph is then processed by a single Graph Convolutional Neural Network (GCN) encoder that learns joint representations capturing both semantic and spatial proximity.
-In the MultiGraph formulation, we construct two independent graphs over the same node set $\mathcal{V}$: a semantic graph $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E}_{s})$, with edge weights:
-$w_{ij}^{(s)}=\cos(\mathbf{h}_{i},\mathbf{h}_{j}),$
-and a geographic graph $\mathcal{G}_{g}=(\mathcal{V},\mathcal{E}_{g})$, with edge weights:
-$w_{ij}^{(g)}=\frac{1}{1+d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})},$
-where $\mathbf{p}_{i}=(\phi_{i},\lambda_{i})$ denotes the latitude–longitude coordinates of node $i$ in degrees under EPSG:4326, and $d_{\mathrm{hav}}$ is the great-circle (Haversine) distance given by:
+Our approach leverages a collection of geotagged social media posts, where each post (node) $i$ is represented by a semantic embedding $\mathbf{h}_{i}$, derived from its textual content, and a geographic coordinate $\mathbf{l}_{i}$.
+Semantic representations are obtained using Sentence-BERT (SBERT), which fine-tunes transformer-based language models in a Siamese architecture to directly optimize sentence-level similarity.
+SBERT has been shown to effectively capture semantic proximity in short, informal texts such as tweets [Reimers2019].
+To enable joint reasoning over heterogeneous data types, all modalities are embedded into graph-based representations that serve as a common computational substrate.
+As already mentioned, graphs provide a flexible framework for modeling local neighborhood structure and relational dependencies induced by semantic similarity and geographic proximity.
+While this study focuses on semantic and geographic signals, the methodology is formulated at the graph level and is therefore modality-agnostic, assuming additional information sources can be easyly encoded as node features or relational graphs.
+We explore two complementary graph construction strategies: a unified MonoGraph formulation and a decoupled MultiGraph formulation.
+Figure 1: Layout of the two models: MonoGraph (a), and MultiGraph (b)
+In the MonoGraph formulation, we construct a single heterogeneous graph $\mathcal{G}_{u}=(\mathcal{V},\mathcal{E}_{u})$ over social media posts, where each node $i\in\mathcal{V}$ corresponds to a post represented by a semantic embedding $\mathbf{h}_{i}$ and geographic coordinates $\mathbf{p}_{i}=(\phi_{i},\lambda_{i})$.
+The graph captures semantic and geographic information through two relation types defined over a shared node set.
+Semantic similarity between nodes is measured using cosine similarity, defining a semantic edge type with weights
+$w_{ij}^{(s)}=\cos(\mathbf{h}_{i},\mathbf{h}_{j}).$
+Geographic proximity is measured using the Haversine distance
 $d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})=2R\arcsin\!\left(\sqrt{\sin^{2}\!\left(\frac{\phi_{j}-\phi_{i}}{2}\right)+\cos(\phi_{i})\cos(\phi_{j})\sin^{2}\!\left(\frac{\lambda_{j}-\lambda_{i}}{2}\right)}\right),$
-where $R$ is the Earth’s radius (e.g., $R\approx 6371\,\mathrm{km}$), and all angular quantities $\phi$ and $\lambda$ are expressed in radians before computation.
-Similarly, this graph is processed with a dedicated GCN network as well.
-Figure 1: Layout of the two models: MonoGraph (a), and MultiGraph (b)
+where $R\approx 6371\,\mathrm{km}$ denotes the Earth’s radius. Geographic proximity defines a geographic edge type with weights
+$w_{ij}^{(g)}=\frac{1}{1+d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})}.$
+For each node $i$, a $k$-nearest-neighbor set is determined based on semantic similarity.
+Both semantic and geographic edge types are instantiated over this shared neighborhood, without introducing explicit modality weighting.
+Graph sparsity is controlled solely through $k$-nearest neighbor selection.
+The resulting heterogeneous graph is processed by a heterogeneous GCN encoder composed of relation-specific graph convolution operators.
+At each layer, messages from different relations are aggregated via summation, enabling joint propagation of semantic similarity and spatial proximity within a unified message-passing framework.
+In the MultiGraph formulation, semantic and geographic relations are instantiated as two independent graphs over the same node set $\mathcal{V}$, allowing modality-specific neighborhood structures to be modeled separately.
+The semantic graph $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E}_{s})$ connects each node to its top-$k$ nearest neighbors based on cosine similarity of semantic embeddings, using the semantic edge weights $w_{ij}^{(s)}$ defined above.
+Graph sparsity is controlled solely through $k$-nearest-neighbor selection, with no global similarity threshold.
+In parallel, the geographic graph $\mathcal{G}_{g}=(\mathcal{V},\mathcal{E}_{g})$ is constructed by connecting each node to its $k$ geographically nearest neighbors based on great-circle distance, with geographic edge weights $w_{ij}^{(g)}$ computed using the Haversine distance defined above.
+No additional thresholds or modality-specific reweighting are introduced.
+Each graph is processed by a dedicated GCN encoder, enabling semantic and geographic neighborhood information to be learned independently prior to fusion.
 3.1 Network Architectures
@@
 1.
-We employ Graph Convolutional Networks (GCNs) to model local context and neighborhood interactions across semantic and geographic dimensions in both architectures.
-GCNs operate via a localized message-passing scheme, updating node features according to:
+We employ Graph Convolutional Networks (GCNs) to model local neighborhood interactions induced by semantic and geographic graph structures.
+GCNs operate via a localized message-passing scheme, updating node features according to
 $\mathbf{X}^{(l+1)}=\sigma\left(\hat{A}\mathbf{X}^{(l)}\mathbf{W}^{(l)}\right),$
 where $\hat{A}$ is the normalized adjacency matrix, $\mathbf{X}^{(l)}$ denotes the feature matrix at layer $l$, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\sigma$ is a non-linear activation function such as ReLU.
-Our implementation uses two GCN layers with ReLU activations.
-In the MultiGraph setup, dropout and layer normalization are applied between layers to enhance generalization and training stability.
-The MonoGraph architecture processes the unified graph with a shared GCN encoder to learn joint node embeddings.
-Conversely, the MultiGraph architecture builds two separate GCN branches, one for semantic and one for geographic graphs, producing semantic embeddings $\mathbf{x}_{i}^{(s)}$ and geographic embeddings $\mathbf{x}_{i}^{(g)}$ for each node $i$.
-These embeddings are then fused through a learnable multi-head attention mechanism.
-In this mechanism, the geographic embeddings serve as the queries $(Q)$, while the semantic embeddings act as both the keys $(K)$ and values $(V)$, enabling the model to selectively attend to relevant semantic features conditioned on geographic context:
-$\mathbf{Z}_{i}=\text{Multi-Head}\big(\mathbf{Q}=\mathbf{x}_{i}^{(g)},\mathbf{K}=\mathbf{x}_{i}^{(s)},\mathbf{V}=\mathbf{x}_{i}^{(s)}\big),\quad\mathbf{z}_{i}=\frac{1}{h}\sum_{j=1}^{h}\mathbf{Z}_{i}^{(j)},$
-where the multi-head attention layer first linearly projects $(Q,K,V)$ into $h$ subspaces (attention heads) of dimension $d_{k}=\frac{\text{embed\_dim}}{h}$, then computes scaled dot-product attention for each head as
-$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
-and finally concatenates the outputs from all attention heads into a single vector, which is then linearly transformed to produce the fused embedding.
-Here, $h$ denotes the number of attention heads.
-This fusion mechanism enables adaptive balancing of semantic and spatial information, selectively modulating the influence of nodes and edges to optimize the overall training objective.
+In both architectures, we use two GCN layers with ReLU activations.
+In the MultiGraph setup, dropout and layer normalization are applied between layers to improve generalization and training stability.
+In the MonoGraph architecture, the unified graph is processed by a single shared GCN encoder, yielding joint node embeddings that integrate semantic similarity and geographic proximity through shared message passing.
+In contrast, the MultiGraph architecture employs two separate GCN branches, one operating on the semantic graph and one on the geographic graph, producing semantic embeddings $\mathbf{x}_{i}^{(s)}$ and geographic embeddings $\mathbf{x}_{i}^{(g)}$ for each node $i$.
+These modality-specific embeddings are subsequently fused through a learnable multi-head cross-attention mechanism.
+In this formulation, geographic embeddings serve as the queries $(Q)$, while semantic embeddings act as both the keys $(K)$ and values $(V)$.
+Since attention is applied independently at the node level rather than over sequences, this mechanism functions as a geo-conditioned feature gating operation, allowing the model to selectively reweight semantic feature dimensions based on geographic context rather than modeling interactions across nodes.
+This asymmetric design is intentional: geographic context provides a stable, low-dimensional conditioning signal that modulates which semantic features are emphasized, rather than being transformed itself.
+Treating geographic embeddings as queries allows the model to adaptively select and reweight semantic dimensions that are most relevant for a given spatial context, while preserving the expressive capacity of the semantic representations.
+Formally, the fused representation is obtained as:
+$\mathbf{z}_{i}=\mathrm{Attn}\bigl(\mathbf{x}_{i}^{(g)},\mathbf{x}_{i}^{(s)},\mathbf{x}_{i}^{(s)}\bigr),$
+where $\mathrm{Attn}(\cdot)$ denotes multi-head scaled dot-product attention.
+The multi-head attention layer first linearly projects $(Q,K,V)$ into $h$ subspaces (attention heads) of dimension $d_{k}=\frac{\text{embed\_dim}}{h}$.
+For each head, scaled dot-product attention is computed as:
+$\mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{softmax}\!\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
+and the outputs of all heads are concatenated and linearly transformed to produce the final fused embedding.
 3.2 Loss Design and Training Objective
-To guide the learning of node representations, we use a composite loss that combines contrastive, coherence, and alignment objectives.
-The contrastive loss helps preserve local semantic structure.
-The coherence loss promotes cluster structure.
-The alignment loss further stabilizes training by encouraging each node to stay close to the centroid of its assigned cluster.
-By jointly optimizing these losses, the model is able to learn embeddings that capture both local neighborhood structure and global cluster organization, while also maintaining alignment with cluster prototypes.
-Clustering is performed using the K-means algorithm with a fixed random seed and fixed cluster numbers to ensure consistency across training epochs, although any other clustering algorithm, differentiable or not, can be used without modifying the core framework.
-Coherence Loss: The coherence loss consists of two components: intra-cluster similarity, which ensures topics are internally consistent, and inter-cluster dissimilarity, which discourages overlap between topics.
-Intra-cluster Similarity: The intra-cluster similarity measures how similar the embeddings of nodes within the same cluster are to each other.
-Intuitively, nodes that belong to the same community should have similar representations, as they share common structural or semantic properties.
-This similarity is commonly quantified using normalized pairwise cosine similarity, which has been shown to be an effective measure for assessing the compactness of clusters, particularly for text and high-dimensional data [ghosh2006].
-The cosine similarity captures the angular distance between embeddings, providing a metric that is robust to differences in magnitude while emphasizing the relative similarity of node features:
-$\mu_{\text{intra}}=\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)=C(j)\}}}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}$
-where $K$ is the number of clusters, and $C_{k}$ denotes the set of nodes in cluster $k$.
-The denominator normalizes the similarity by the number of node pairs in each cluster.
-Cosine similarity is computed as:
-$\cos(\mathbf{z}_{i},\mathbf{z}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|_{2}\|\mathbf{z}_{j}\|_{2}}$
-Here, $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ represent the feature embeddings of nodes $i$ and $j$, respectively, and are normalized by their L2 norms to ensure that only the direction (rather than magnitude) of the vectors influences the similarity measure.
-Maximizing intra-cluster similarity encourages the model to learn node embeddings that are close to one another within the same cluster.
-This promotes compactness and internal consistency in the embedding space, improving the semantic or spatial cohesion of discovered clusters.
-In this context, higher values of intra-cluster similarity, typically close to 1, indicate tight and cohesive clusters, while values near 0 or -1 suggest poorly defined clusters.
-Inter-cluster Similarity: The inter-cluster similarity captures how similar node embeddings are across different clusters.
-Ideally, embeddings from separate clusters should be dissimilar.
-We define this as:
-$\mu_{\text{inter}}=\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})\cdot\mathbf{1}_{\{C(i)\neq C(j)\}}}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}$
-Here, $P_{\text{inter}}$ denotes all node pairs from distinct clusters.
-High inter-cluster similarity values would indicate insufficient separation between clusters, while values close to -1 reflect distinct and well-separated clusters.
-So, ideally we aim for low inter-cluster similarity, with values closer to -1 indicating optimal separation between clusters.
-To jointly optimize intra-cluster compactness and inter-cluster separation, we define the coherence loss:
-$\mathcal{L}_{\text{coherence}}=-\left(\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}-\lambda\cdot\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}\right)$
-where $\lambda$ is a hyperparameter that controls the trade-off between maximizing intra-cluster similarity and minimizing inter-cluster similarity.
-This loss function combines both clustering principles into a single differentiable objective.
-The negative sign transforms the coherence score into a minimization problem, consistent with standard optimization procedures.
-Minimizing $\mathcal{L}_{\text{coherence}}$ leads to embedding spaces where nodes within the same cluster are tightly grouped, while nodes from different clusters are well-separated, enhancing both interpretability and performance of downstream clustering tasks.
-Contrastive Loss: To preserve local semantic or structural neighborhoods, we use a contrastive loss that pulls embeddings of similar nodes closer and pushes dissimilar ones apart.
-For a batch of positive and negative node pairs, the contrastive objective is defined as:
-$\mathcal{L}_{\text{contrast}}=-\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{(i,k)\in\mathcal{N}_{i}}\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
-where $\mathcal{P}$ denotes the set of positive node pairs (e.g., semantically or spatially close), $\mathcal{N}_{i}$ includes both positives and negatives for node $i$, and $\tau$ is a temperature hyperparameter controlling the sharpness of the distribution.
-Contrastive loss enforces local consistency in the learned embeddings, ensuring that nodes with similar contexts (semantic or spatial) are embedded close together.
-This regularizes the representation space and improves generalization, especially in heterogeneous graphs.
-Alignment Loss: The alignment loss encourages node embeddings to remain close to the centroid of their assigned cluster.
-This constraint stabilizes training and reduces noise by enforcing tighter adherence to cluster prototypes.
-The loss is defined as:
-$\mathcal{L}_{\text{align}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{c}_{C(i)}\right\|_{2}^{2}$
-where $\mathbf{z}_{i}$ is the embedding of node $i$, and $\mathbf{c}_{C(i)}$ is the centroid of cluster $C(i)$, computed as the mean embedding of all nodes in that cluster.
-This loss helps to maintain cluster cohesion by anchoring each node close to its corresponding cluster center.
-When combined with coherence and contrastive losses, alignment loss enhances the structural integrity of the latent space, especially in high-noise environments or early training stages.
-Total Loss Objective: To effectively learn cluster-aware and structure-preserving embeddings, we combine all three objectives into a unified loss function. The total loss is defined as:
-$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{contrast}}+\beta\cdot\mathcal{L}_{\text{coherence}}+\gamma\cdot\mathcal{L}_{\text{align}}$
-where $\alpha$, $\beta$, and $\gamma$ are weighting coefficients that control the relative contribution of each component.
-This composite formulation allows the model to jointly optimize for local consistency, global cluster structure, and prototype alignment.
+To guide the learning of node representations, we optimize a composite objective that combines contrastive, coherence, and alignment losses.
+The contrastive loss preserves local semantic neighborhood structure induced by the graph, while the coherence and alignment losses introduce global structural regularization by encouraging compact and well-separated clusters in the embedding space.
+Joint optimization of these objectives enables the model to learn representations that capture both local relational consistency and global cluster organization.
+During training, clustering is incorporated to provide coarse-grained global guidance that complements the local nature of contrastive learning.
+At each training epoch, node embeddings are clustered using the K-means algorithm with a fixed number of clusters ($k_{\text{means}}=15$) and a fixed random seed ($42$) to ensure consistency across epochs.
+The resulting cluster assignments are used exclusively to compute the coherence and alignment losses and act as dynamic pseudo-labels, while the contrastive loss operates solely on local graph neighborhoods.
+Since clustering is non-differentiable and performed outside the computational graph, it does not impose hard constraints on the embedding geometry.
+For evaluation and qualitative analysis, clustering is performed post hoc on the final learned embeddings using Spectral Clustering, as described in Section 3.3.
+Coherence Loss: The coherence loss consists of two complementary components: intra-cluster similarity, which promotes internal consistency within clusters, and inter-cluster similarity, which discourages overlap between different clusters.
+The intra-cluster similarity measures how similar the embeddings of nodes within the same cluster are to each other:
+$\mu_{\text{intra}}=\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}$,
+where $K$ is the number of clusters and $C_{k}$ denotes the set of nodes assigned to cluster $k$. Cosine similarity is computed as
+$\cos(\mathbf{z}_{i},\mathbf{z}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|_{2}\|\mathbf{z}_{j}\|_{2}}.$
+Maximizing intra-cluster similarity encourages compact and internally consistent clusters in the learned embedding space.
+The inter-cluster similarity captures the similarity between embeddings belonging to different clusters:
+$\mu_{\text{inter}}=\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}},$
+where $P_{\text{inter}}$ denotes all node pairs drawn from distinct clusters.
+Lower values indicate better separation between clusters.
+To jointly encourage intra-cluster compactness and inter-cluster separation, the coherence loss is defined as:
+$\mathcal{L}_{\text{coherence}}=-\left(\frac{\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i,j\in C_{k}\\ i\neq j\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{k=1}^{K}\binom{|C_{k}|}{2}}-\lambda\cdot\frac{\sum_{\begin{subarray}{c}i,j\in P_{\text{inter}}\\ C(i)\neq C(j)\end{subarray}}\cos(\mathbf{z}_{i},\mathbf{z}_{j})}{\sum_{i\neq j}\mathbf{1}_{\{C(i)\neq C(j)\}}}\right),$
+where $\lambda$ controls the trade-off between the two terms.
+Contrastive Loss: To preserve local semantic or structural neighborhoods, we employ a contrastive loss that pulls embeddings of similar nodes closer while pushing dissimilar ones apart.
+Positive pairs are constructed from the semantic graph: for each node $i$, its positive set consists of its top-$k$ semantic neighbors connected to $i$ by edges in the semantic graph, while all other nodes within the same mini-batch act as implicit negatives.
+The contrastive objective is defined as:
+$\mathcal{L}_{\text{contrast}}=-\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{(i,k)\in\mathcal{N}_{i}}\exp(\cos(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)},$
+where $\mathcal{P}$ denotes the set of positive node pairs, $\mathcal{N}_{i}$ contains both positive and negative samples for node $i$, and $\tau$ is a temperature parameter controlling the sharpness of the similarity distribution.
+Unless otherwise stated, we use a fixed temperature parameter $\tau=0.5$ during training, as identified by the sensitivity analysis in Section 6.
+Alignment Loss: The alignment loss encourages node embeddings to remain close to the centroid of their assigned cluster:
+$\mathcal{L}_{\text{align}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{z}_{i}-\mathbf{c}_{C(i)}\right\|_{2}^{2},$
+where $\mathbf{c}_{C(i)}$ denotes the centroid of cluster $C(i)$, computed as the mean embedding of all nodes assigned to that cluster.
+Total Loss Objective: The final training objective combines all components as:
+$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{contrast}}+\beta\cdot\mathcal{L}_{\text{coherence}}+\gamma\cdot\mathcal{L}_{\text{align}},$
+where $\alpha$, $\beta$, and $\gamma$ control the relative contribution of each loss term.
+Unless otherwise stated, we use the baseline configuration $\alpha=0.8$, $\beta=0.2$, $\gamma=0.1$, $k_{\text{means}}=15$, $seed=42$, and $\tau=0.5$, during training, as identified by the sensitivity analysis in this section.
 3.3 Embedded Features Clustering
@@
 This method is particularly effective for high-dimensional and non-convex data distributions, such as those found in graph-based embeddings [von2007].
-Likewise, we fix the number of clusters $K=10$ to stabilize the process and reduce variability, ensuring consistent topic identification across runs.
-This choice also enhances reproducibility by maintaining the same clustering configuration.
-4 Validation
-A comprehensive validation was conducted to rigorously assess the performance and generalization capability of the proposed methodology.
-We employed a multi-faceted experimental procedure for this purpose.
-First, we evaluated the models’ ability to capture coherent patterns and embed the social media data into clusters that are cohesive and distinct, corresponding to high intra-cluster similarity and low inter-cluster similarity.
-Next, we benchmarked their performance against the most comparable method identified in the literature, as well as a sequential baseline, across four diverse datasets.
-Finally, the utility of the proposed models was demonstrated through a case study on the 2017 Hurricane Harvey in the United States, in which we evaluated the interpretability and the spatial autocorrelation patterns found in the relevant topics, which are proven to be significant for disaster management applications.
-4.1 Experimental Setup
-Datasets: As the models were designed specifically for the analysis of social media posts in disaster situations, we validated our proposed models using four tweet datasets related to distinct real-world disaster scenarios: (1) the 2014 Napa earthquake [brocher2015], which occurred on August 24, 2014.
-The geographical bounds focus on the Napa Valley region in California, covering 4,287 km².
-(2) Hurricane Harvey in 2017 [noaa2017], a Category 4 hurricane that impacted the southeastern Texas region from August 25 to September 7, 2017.
-The dataset reflects a large spatial extent of 26,736 km², particularly around Houston.
-(3) The 2021 Ahr Valley flood [koks2021], caused by heavy rainfall, which took place in July 2021.
-The dataset covers the 1,256 km² area of the Ahr Valley in Germany.
-(4) The 2023 Chile wildfires [cordero2024], which affected central and southern Chile and were most intense in February 2023.
-The dataset spans a vast area of 546,905 km².
-All the datasets were collected using Twitter’s former v1.1 streaming and recent search API endpoints, while filtering by the bounding box Polygon (Geographic Bounds), and the posting time (Time Frame), as depicted in Table 1.
-As a common practice in natural language processing (NLP), we preprocessed the raw text data (tweets) before performing further analysis.
-This preprocessing is done to ensure consistency, remove noise, and improve the quality of the data.
-Specifically, we removed URLs using a regular expression and converted the text to lowercase for keyword extraction.
-HTML entities were unescaped, Mentions (i.e., @references) were removed, and extra whitespace was stripped.
-Non-alphanumeric characters were removed, followed by the elimination of numbers.
-For BERT-specific cleaning, Mentions were replaced with @user, URLs were converted to http, and newline characters were removed.
-These steps ensure that the tweets are appropriately formatted for further analysis and model training.
+We fix the number of clusters ($k_{\text{spect}}=10$) for post hoc analysis to stabilize the clustering process and reduce variability, ensuring consistent topic identification across runs.
+This choice is made independently of the training-time clustering used for representation learning and reflects a trade-off between topic granularity and interpretability commonly adopted in topic modeling evaluation.
+Fixing $k_{\text{spect}}$ also enhances reproducibility by maintaining a consistent clustering configuration.
+4 Experimental Setup
+This section describes the datasets, preprocessing steps, model configuration, and training protocol used throughout the experimental evaluation.
+All design choices are reported once here to ensure consistency and reproducibility across subsequent analyses, including clustering evaluation and Topic Quality (TQ) benchmarking.
+4.1 Datasets
+As the models are designed specifically for the analysis of social media posts in disaster situations, we validated our proposed methods using four tweet datasets corresponding to distinct real-world disaster events: (1) the 2014 Napa earthquake [brocher2015], which occurred on August 24, 2014.
+The geographic bounds focus on the Napa Valley region in California, covering 4,287 km2; (2) Hurricane Harvey in 2017 [noaa2017], a Category 4 hurricane impacting southeastern Texas between August 25 and September 7, 2017, covering a spatial extent of 26,736 km2; (3) the 2021 Ahr Valley flood [koks2021], which occurred in July 2021 and affected a 1,256 km2 region in Germany; and (4) the 2023 Chile wildfires [cordero2024], which impacted central and southern Chile in February 2023 and span a large area of 546,905 km2.
+All datasets were collected using Twitter’s former v1.1 streaming and recent search API endpoints, filtered by geographic bounding polygons and event-specific time frames  [^1] The data supporting the findings of this study are available from the authors upon reasonable request, subject to platform terms and conditions.
+.
+Table 1 summarizes the key properties of each dataset.
 Table 1: Properties of the datasets used for validation.
-The time frame and geographic bounds were used for filtering.
-In addition, we computed the number of posts within each dataset and the size of the cover area.
 Dataset Posts Time Frame Geographic Bounds Cover Area
-Ahr Valley floods 11,177 2021-07-01 – 2021-07-31 (50.01, 6.00, 50.75, 8.00) 1,256 km²
-Hurricane Harvey 28,933 2017-08-25 – 2017-09-07 (27.58, –98.77, 30.66, –90.86) 26,736 km²
-Chile wildfires 100,707 2023-02-01 – 2023-02-28 (–55.95, –76.89, –17.34, –65.22) 546,905 km²
-Napa earthquake 90,907 2014-08-24 – 2014-08-25 (37.11, –123.05, 38.98, –121.04) 4,287 km²
-Network Training and Parametrization: The models was trained in an unsupervised manner, without access to ground truth topic or cluster labels, using the compound loss, as detailed in the Section 3.2.
-Input features consisted of SBERT-generated semantic embeddings for textual content and corresponding geolocation coordinates for graph construction.
-These modalities were used to define semantic and geographic edges in both the MonoGraph and MultiGraph formulations.
-Training was performed using the Adam optimizer with an initial learning rate of 1e-3.
-Models were trained for 50 epochs, with early stopping based on the stability of the coherence score over a held-out validation set.
-In the MultiGraph setup, we applied a dropout rate of 0.3 and used two Torch GCN layers per branch.
-The multi-head attention mechanism used for embedding fusion consisted of 4 heads and incorporated layer normalization between fusion stages.
-To efficiently manage memory, avoid over-sized graphs, and enable consistent training across varying dataset sizes, we adopted a strided chunking protocol.
-In specific, each dataset was partitioned into multiple non-overlapping chunks using a fixed stride value.
-Given a stride $s$, each chunk consisted of every $s$-th entry from the original dataset, starting from a different offset.
-Formally, we generated $s$ subsets $\{\text{chunk}_{i}\}_{i=1}^{s}$ such that:
-$\text{chunk}_{i}=\text{Input}_{\text{gdf}}[i::s]$
-This approach preserved the overall diversity and temporal structure of the data while enabling the model to process each chunk independently.
-The protocol ensures that all data points are covered, and total recovery of the original dataset was verified before training.
-Each chunk was used sequentially as input to the model in a looped training cycle.
-Hyperparameter tuning was carried out via a random search over a validation subset of the datasets.
-Ultimately, we set the contrastive, coherence, and alignment loss weights to $\alpha$ = 0.8, $\beta$ = 0.2, and $\gamma$ = 0.1, respectively.
-These choices are based on a combination of domain knowledge, empirical tuning, and the relative importance of each loss component for the task at hand.
-For the contrastive loss, we set the temperature parameter $\tau$ = 0.5.
-All experiments were executed on an NVIDIA H100 GPU.
+Ahr Valley floods 11,177 2021-07-01 – 2021-07-31 (50.01, 6.00, 50.75, 8.00) 1,256 km2
+Hurricane Harvey 28,933 2017-08-25 – 2017-09-07 (27.58, –98.77, 30.66, –90.86) 26,736 km2
+Chile wildfires 100,707 2023-02-01 – 2023-02-28 (–55.95, –76.89, –17.34, –65.22) 546,905 km2
+Napa earthquake 90,907 2014-08-24 – 2014-08-25 (37.11, –123.05, 38.98, –121.04) 4,287 km2
+4.2 Preprocessing
+As is standard practice in natural language processing, the raw tweet text was preprocessed prior to analysis to ensure consistency and reduce noise.
+URLs were removed using regular expressions and all text was lowercased.
+HTML entities were unescaped, user mentions were removed, and excess whitespace was stripped.
+Non-alphanumeric characters and numeric tokens were eliminated.
+For embedding generation, we follow the input preprocessing conventions recommended for Sentence-BERT, replacing mentions with the token @user, mapping URLs to http, and removing newline characters.
+These steps ensure compatibility with the pre-trained language models and consistency across datasets.
+4.3 Model Configuration and Training
+All models were trained in an unsupervised manner without access to ground-truth topic or cluster labels, using the compound objective described in Section 3.2.
+Input features consisted of SBERT-generated semantic embeddings and associated geolocation coordinates, which were used to construct semantic and geographic graphs in both the MonoGraph and MultiGraph formulations.
+Training was performed using the Adam optimizer with an initial learning rate of $1\times 10^{-3}$.
+Models were trained for 50 epochs with early stopping based on coherence stability on a held-out validation subset.
+In the MultiGraph configuration, each branch employed two GCN layers with a dropout rate of 0.3.
+Fusion was performed via multi-head attention with four heads, and layer normalization was applied between fusion stages to stabilize training.
+4.4 Scalability via Strided Chunking
+To scale graph construction to large corpora, we adopt strided chunking during preprocessing. Given a stride $s$, the dataset is partitioned into $s$ partially overlapping subsets defined as
+$\text{chunk}_{i}=\text{Input}_{\mathrm{gdf}}[i::s],\quad i=1,\dots,s.$
+This construction ensures that each node appears in exactly one position per stride cycle while nodes near chunk boundaries are shared across multiple chunks, preserving local semantic and geographic neighborhoods.
+Chunk overlap prevents artificial fragmentation of neighborhoods that could otherwise arise from independent graph construction.
+As a result, node embeddings remain consistent across chunks, and striding affects computational efficiency only, not the learned representation or clustering behavior.
+Full data coverage was verified prior to training, and all chunks were processed sequentially under identical optimization settings.
+Empirically, varying the stride within reasonable bounds did not result in meaningful changes in the embedding geometry or TQ, confirming that strided chunking functions strictly as a scalability mechanism and does not bias embedding geometry or topic structure.
+Stride values are selected based on memory constraints rather than dataset-specific tuning, and do not affect neighborhood structure or learned representations.
+4.5 Baseline Configuration
+The primary baseline for all experiments is the sequential semantic workflow introduced in this study.
+It operates exclusively on semantic embeddings and does not incorporate graph-based modeling, geographic information, multimodal fusion, or contrastive learning.
+This configuration serves as a minimal reference point and enables direct attribution of performance differences to the proposed graph construction, semantic–geographic integration, and contrastive objectives.
+In addition, we include the JSTTS multimodal pipeline proposed by Hanny et al.
+[hanny2025] as an external reference method.
+Although temporal and sentiment features are part of the original JSTTS formulation, these modalities are intentionally excluded from all primary experiments in this work.
+This design choice applies uniformly across all evaluated models and is motivated by the goal of isolating the methodological contribution of semantic–geographic graph modeling and contrastive learning.
+Incorporating auxiliary temporal or affective signals would alter the input space and introduce additional information that is orthogonal to the methodological focus of this study, thereby confounding attribution of performance differences.
+Nonetheless, for completeness and transparency, the behavior of JSTTS under its original full-modality configuration is reported separately in the evaluation section.
+These results are provided for contextual reference only and are not used for direct comparison with the proposed models.
+To ensure comparability with prior work, we use the same datasets and embedding backbones as in [hanny2025] across all experiments, including clustering and topic quality evaluation.
+Specifically, the English all-MiniLM-L12-v2 model is used for the Napa earthquake and Hurricane Harvey datasets, while the multilingual distiluse-base-multilingual-cased-v1 model is used for the Ahr Valley floods and Chile wildfires datasets.
+All models are evaluated using identical preprocessing pipelines, clustering procedures, and keyword extraction methods.
+5 Evaluation
+This section evaluates the proposed models from two complementary perspectives.
+First, we assess their ability to produce semantically coherent and well-separated clusters in the embedding space.
+Second, we benchmark topic modeling performance using TQ against a sequential baseline and an established multimodal pipeline.
+Together, these evaluations quantify both representation quality and downstream topic modeling effectiveness across diverse disaster scenarios.
 Table 2: Clustering evaluation results across different datasets using the two models against the input referenced data
@@
 5.1 Clustering Capability Evaluation
-To evaluate the models’ capacity for generating semantically coherent representations, we conducted clustering assessments on both the original SBERT input embeddings and the output embedded features produced by each model.
-Following the methodology outlined in Section 3.2, we computed two key metrics: intra-cluster similarity, reflecting semantic cohesion within clusters, and inter-cluster similarity, indicating the degree of separation between them.
-Ideally, effective clustering is characterized by high intra-cluster cohesion and low inter-cluster similarity.
-Table 2 summarizes the results across the four disaster-specific datasets.
-The input embeddings serve as a baseline and consistently exhibit moderate cohesion but relatively poor separation, suggesting that the raw SBERT representations lack strong semantic structure.
-Both MonoGraph and MultiGraph yield substantial improvements.
-Across all datasets, inter-cluster similarity drops significantly, while intra-cluster cohesion improves.
-MultiGraph achieves the highest cohesion values on all datasets and attains the lowest (i.e., most negative) inter-cluster similarity scores for the Chile wildfires and Napa earthquake datasets, suggesting well-isolated and internally consistent clusters.
-On the Ahr Valley floods and Hurricane Harvey datasets, MonoGraph shows better (lower) inter-cluster similarity values, although its intra-cluster scores are slightly lower than MultiGraph’s, yet still markedly better than the baseline.
-These trends are further reinforced by the visualizations in Fig.
+To evaluate the models’ capacity for generating semantically coherent representations, we conducted clustering assessments on both the original input embeddings and the output embedded features produced by each model.
+Following the methodology outlined in Section 3.2, we computed two complementary metrics: intra-cluster similarity, reflecting semantic cohesion within clusters, and inter-cluster similarity, indicating the degree of separation between clusters.
+Effective clustering is characterized by high intra-cluster cohesion and low inter-cluster similarity.
+Table 2 summarizes the results across four disaster-specific datasets.
+The raw input embeddings serve as a reference baseline and consistently exhibit moderate intra-cluster cohesion but relatively poor inter-cluster separation, indicating that semantic structure alone is insufficient for robust clustering.
+Both MonoGraph and MultiGraph substantially improve clustering behavior across all datasets.
+Inter-cluster similarity decreases markedly, while intra-cluster cohesion increases, indicating tighter and better-separated clusters.
+MultiGraph achieves the highest intra-cluster cohesion across all datasets and yields the lowest (i.e., most negative) inter-cluster similarity for the Chile wildfires and Napa earthquake datasets, suggesting particularly well-isolated and internally consistent clusters.
+For the Ahr Valley floods and Hurricane Harvey datasets, MonoGraph attains slightly lower inter-cluster similarity values, although its intra-cluster cohesion remains marginally below that of MultiGraph while still substantially outperforming the baseline.
+These quantitative findings are reinforced by the visualizations in Fig.
 2.
-The input similarity matrix (top-left) exhibit less well-defined structure, whereas the output matrices (middle-left and bottom-left) display sharper contrast and more structured groupings.
-Likewise, the 3D cluster plots (right) reveal a transition from scattered distributions in the input space to more compact and semantically coherent clusters after model transformation.
-Low inter-cluster similarity values indicate strong dissimilarity between clusters, reflecting effective cluster separation.
-In summary, both models, especially MultiGraph, demonstrate a strong ability to reorganize the input features space in ways that enhance clustering quality.
-This improvement is crucial for downstream tasks such as topic modeling, trend detection, and spatial-semantic analysis.
-4.3 Benchmark and Baseline Comparison
-The objective of this evaluation was to quantitatively assess the models’ performance and their capability to generate high-quality clusters across diverse disaster scenarios.
-For benchmarking purposes, we compared our proposed models against the multimodal pipeline presented by Hanny et al.
-[hanny2025], as well as the baseline sequential workflow introduced in this study.
-To rule out the effect of time and sentiment features in the multimodal pipeline, we set the sentiment and time weights in the pipeline to zero during the evaluation.
-To ensure fair comparison, we used the same exact datasets as used in [hanny2025].
-For embedding generation, we utilized the English-only all-MiniLM-L12-v2 SBERT model for the 2014 Napa earthquake and 2017 Hurricane Harvey datasets.
-Conversely, for the 2021 Ahr Valley flood and 2023 Chile wildfires datasets, we adopted the multilingual distiluse-base-multilingual-cased-v1 model, consistent with the approach in [hanny2025].
-To ensure a congruent comparison across models, we chose Topic Quality (TQ), a widely used metric in the community.
-TQ effectively evaluates the coherence and relevance of topics, making it a reliable measure for assessing the quality of topic models.
-It is computed based on the top 10 keywords per topic and defined as the product of Topic Coherence (TC) and Topic Diversity (TD).
-We employed a TF-IDF-based keyword extraction method: For each cluster, the corresponding subset of documents was aggregated, and TF-IDF scores were calculated over this localized corpus.
-The top 10 highest-scoring terms were selected as representative keywords.
-To compute the TQ score, we first constructed a document-term matrix using CountVectorizer, then estimated word and word-pair frequencies across the full corpus.
-For each topic, pairwise Positive Pointwise Mutual Information (PPMI) between all keyword pairs was calculated and normalized by the negative log joint probability to obtain the normalized variant (NPPMI) [bouma2009] [jurafsky2023].
-Formally, the normalized PPMI between two words $w_{i}$ and $w_{j}$ is given by
-$\text{NPPMI}(w_{i},w_{j})=\frac{\max(0,\log\frac{P(w_{i},w_{j})}{P(w_{i})P(w_{j})})}{-\log P(w_{i},w_{j})},$
-where probabilities are estimated from co-occurrence counts over the corpus.
-The coherence of a topic is the average NPPMI over all $\binom{10}{2}$ word pairs in the topic, and the overall Topic Coherence is the mean across all valid topics.
-Topic Diversity quantifies how distinct the discovered topics are by measuring the proportion of unique keywords across all topics.
-It is calculated as
+The input similarity matrix exhibits diffuse structure, whereas the learned representations display sharper contrast and clearer block patterns.
+Likewise, the 3D visualizations reveal a transition from scattered point clouds to compact and well- separated clusters.
+Overall, both models, and especially MultiGraph, demonstrate a strong ability to reorganize the embedding space in a way that enhances clustering quality, which is critical for downstream topic modeling and spatial-semantic analysis.
+5.2 Topic Quality Benchmarking
+This section evaluates the semantic quality of the topics discovered by the proposed models under controlled and comparable conditions.
+We first describe the evaluation protocol and baseline configurations used for TQ assessment, then formally define the TQ metric, and finally present quantitative results across all datasets.
+5.2.1 Evaluation Protocol
+Following the baseline definitions established in Section 4.5, TQ evaluation is conducted under controlled and directly comparable conditions.
+All models are evaluated using semantic embeddings as the common input modality.
+Both the multimodal models additionally incorporate geographic information, while the BERTopic baseline operates solely on semantic representations, reflecting its original formulation.
+In addition to the primary benchmarking experiments, we report TQ scores for the JSTTS pipeline proposed by Hanny et al.
+[hanny2025] in its original intended configuration, with all available modalities, including semantic, spatial, temporal, and sentiment features, enabled (reported in Table 3).
+As already mentioned in Section 4.5, introducing auxiliary information that is orthogonal to the methodological focus of this study.
+As a result, these results are not treated as directly comparable benchmarks, but are reported for transparency and contextual reference, illustrating the relative scale of TQ achieved by modular multimodal pipelines.
+5.2.2 Topic Quality Metric
+Topic quality is a widely adopted metric for evaluating topic models, as it jointly captures semantic coherence and lexical diversity.
+It is defined as the product of Topic Coherence (TC) and Topic Diversity (TD), computed using the top-$k$ keywords per topic (with $k_{\text{means}}=15$ in all experiments).
+To extract representative keywords, documents assigned to each cluster are aggregated and a TF–IDF vectorizer is applied locally.
+The top-$k$ terms with the highest TF–IDF scores are selected as topic descriptors.
+Topic Coherence is computed using normalized Positive Pointwise Mutual Information (NPPMI) [bouma2009, jurafsky2023].
+Given two words $w_{i}$ and $w_{j}$, NPPMI is defined as:
+$\text{NPPMI}(w_{i},w_{j})=\frac{\max\left(0,\log\frac{P(w_{i},w_{j})}{P(w_{i})P(w_{j})}\right)}{-\log P(w_{i},w_{j})},$
+where probabilities are estimated from word co-occurrence counts over the full corpus.
+The coherence of a topic is computed as the average NPPMI over all $\binom{k}{2}$ keyword pairs, and the overall Topic Coherence is obtained by averaging across all valid topics.
+Topic Diversity measures how distinct the discovered topics are by quantifying the proportion of unique keywords across all topics:
 $\text{TD}=\frac{U}{k\times|T|},$
-where $U$ denotes the number of unique words in the top-$k$ keywords across all topics, $k$ is the number of keywords per topic, and $|T|$ is the total number of topics.
-This penalizes repeated keywords across topics and rewards diversity in word usage.
-Finally, Topic Quality is computed as the product of the two metrics:
+where $U$ denotes the number of unique words among the top-$k$ keywords across all topics, and $|T|$ is the total number of topics.
+This formulation penalizes repeated keywords and rewards lexical diversity.
+Finally, TQ is defined as:
 $\text{TQ}=\text{TC}\times\text{TD},$
-which favors models that produce both semantically coherent and lexically diverse clusters. All metrics are computed using the same preprocessing pipeline for consistency across evaluated methods.
+favoring models that produce topics that are both semantically coherent and lexically diverse.
+All components of TQ are computed using the same preprocessing and keyword extraction pipeline for all evaluated methods.
+Table 3: Topic Quality(TQ) achieved by the JSTTS pipeline when all modalities (semantic, spatial, temporal, and sentiment) are enabled, as originally specified.
+       Dataset        JSTTS TQ
+       Ahr Valley Floods        0.165
+       Hurricane Harvey        0.191
+       Chile Wildfires        0.142
+       Napa Earthquake        0.081
 5.3 TQ Results and Analysis
-The Topic Quality (TQ) scores reported in Table 3 clearly demonstrate the effectiveness of our proposed models in generating high-quality topic clusters across diverse disaster-related datasets.
+The TQ scores reported in Table 4 clearly demonstrate the effectiveness of our proposed models in generating high-quality topic clusters across diverse disaster-related datasets.
 The MultiGraph consistently achieves the highest TQ scores in three out of four scenarios, the Ahr Valley floods (0.415), Hurricane Harvey (0.384), and the Napa earthquake (0.195), demonstrating its robustness and adaptability to datasets with varying linguistic, geographic, and topical characteristics.
@@
 In particular, contexts characterized by fewer high-quality labeled signals or more multilingual noise may favor the simpler and more direct MonoGraph architecture, which consistently outperforms baseline methods across all datasets and yields stable, focused clusters despite its relative simplicity.
-Table 3: Topic quality scores (TQ) for outputs from our two models (MonoGraph and MultiGraph), the method by Hanny et al. [hanny2025], and the baseline sequential workflow across four datasets.
+Table 4: TQ for outputs from our two models (MonoGraph and MultiGraph), the method by Hanny et al. [hanny2025], and the baseline sequential workflow across four datasets.
 Dataset MonoGraph MultiGraph Hanny et al. [hanny2025] Sequential Workfollow
@@
 The absence of feedback or interaction between stages appears to hinder its ability to produce globally meaningful topic representations.
-Figure 3: Topic-word distributions for a subset of the Hurricane Harvey dataset, applying MonoGraph (top) and the MultiGraph (bottom) models.
+6 Ablation, Sensitivity, and Stability Analysis
+To provide a systematic understanding of the proposed multimodal graph-based topic modeling framework, we conduct a comprehensive set of ablation, sensitivity, and stability experiments.
+These analyses are designed to: (i) isolate the contribution of individual loss components and architectural choices, and (ii) assess the robustness of the model with respect to hyperparameter settings and stochastic initialization.
+All experiments are conducted on the Hurricane Harvey dataset using both the MonoGraph and MultiGraph variants.
+Performance is evaluated using Topic Quality, which jointly reflects topic coherence and diversity.
+Unless otherwise stated, results reported in this section are averaged over the corresponding experimental settings, and summary statistics are provided in the form of Mean, standard deviation (STD), and 95% Confidence Interval (CI) where applicable.
+Table 5: Loss ablation results based on TQ. ✓ indicates the loss is included, while $\times$ indicates it is excluded.
+   Model    $\mathcal{L}_{c}$    $\mathcal{L}_{h}$    $\mathcal{L}_{a}$    Removed Loss    TQ
+   MonoGraph    ✓    ✓    ✓    None (Full)    0.368
+   MonoGraph    $\times$    ✓    ✓    Contrastive    0.205
+   MonoGraph    ✓    $\times$    ✓    Coherence    0.357
+   MonoGraph    ✓    ✓    $\times$    Alignment    0.353
+   MultiGraph    ✓    ✓    ✓    None (Full)    0.384
+   MultiGraph    $\times$    ✓    ✓    Contrastive    0.218
+   MultiGraph    ✓    $\times$    ✓    Coherence    0.362
+   MultiGraph    ✓    ✓    $\times$    Alignment    0.366
+6.1 Loss Ablation Analysis
+We analyze the contribution of each loss component, contrastive loss ($\mathcal{L}_{c}$), coherence loss ($\mathcal{L}_{h}$), and alignment loss ($\mathcal{L}_{a}$), by removing one term at a time while keeping all other hyperparameters fixed at their baseline configuration.
+Results are summarized in Table 5.
+Removing the contrastive loss leads to a substantial degradation in performance for both model variants, with TQ dropping sharply relative to the full configuration.
+This indicates that contrastive learning provides the primary discriminative signal necessary for effective topic separation.
+In contrast, removing either the coherence or alignment loss results in only moderate reductions in TQ, suggesting that these components act as stabilizing regularizers rather than dominant drivers of clustering quality.
+Overall, the ablation results confirm that the loss formulation is well balanced: contrastive learning establishes topic structure, while coherence and alignment losses refine and stabilize the learned representations.
+Table 6: Sensitivity analysis of loss weights based on Topic TQ.
+     Setting      MonoGraph      MultiGraph
+     $\alpha=0.2$      0.308      0.346
+     $\alpha=0.5$      0.351      0.355
+     $\alpha=0.8$      0.368      0.384
+     $\beta=0.1$      0.344      0.363
+     $\beta=0.2$      0.368      0.384
+     $\beta=0.5$      0.365      0.354
+     $\lambda=0.05$      0.361      0.367
+     $\lambda=0.1$      0.368      0.384
+     $\lambda=0.2$      0.349      0.362
+     Mean      0.354      0.367
+     STD      0.019      0.014
+     95% CI      $\pm$0.013      $\pm$0.009
+Table 7: Sensitivity analysis for contrastive temperature $\tau$ and number of clusters ($k_{\text{means}}=15$) using TQ.
+   Parameter    Value    MonoGraph TQ    MultiGraph TQ
+   $\tau$    0.2    0.314    0.359
+   $\tau$    0.5    0.368    0.384
+   $\tau$    1.0    0.315    0.359
+   $k_{\text{means}}=15$    10    0.341    0.351
+   $k_{\text{means}}=15$    15    0.368    0.384
+   $k_{\text{means}}=15$    20    0.364    0.370
+   Mean    –    0.345    0.368
+   STD    –    0.026    0.017
+   95% CI    –    $\pm$0.021    $\pm$0.014
+Table 8: Stability across random seeds computed over independent runs.
+  Model   Mean TQ   STD   95% CI   Seeds
+  MonoGraph   0.3667   0.0025   $\pm$0.0062   {42,1,2,3}
+  MultiGraph   0.3625   0.0034   $\pm$0.0084   {42,1,2,3}
+6.2 Hyperparameter Sensitivity Analysis
+We assess robustness to hyperparameter selection by independently varying key parameters while keeping all others fixed at their baseline configuration.
+Table 6 summarizes TQ under different loss-weight configurations.
+For MonoGraph, the mean TQ across all tested settings is $0.354$ with a standard deviation of $0.019$, while MultiGraph achieves a mean of $0.367$ with a standard deviation of $0.014$.
+The relatively small dispersion indicates that performance is stable under moderate variations of loss weights, with the baseline configuration consistently yielding the highest TQ.
+Sensitivity with respect to the contrastive temperature $\tau$ and the number of clusters ($k_{\text{means}}=15$) is reported in Table 7.
+TQ exhibits a unimodal trend with respect to $\tau$, where intermediate values provide optimal separation of positive and negative pairs.
+Across all tested values, the standard deviation remains below $0.03$ for both models, indicating limited sensitivity.
+Similarly, performance remains stable across different values of ($k_{\text{means}}$), with peak TQ achieved at the baseline configuration $k_{\text{means}}=15$.
+It is also worth noting that while the absolute topic granularity varies with the choice of $k_{\text{means}}$, the relative performance ranking between MonoGraph and MultiGraph remains stable across the tested range.
+These results demonstrate that the proposed framework is robust to reasonable variations in both optimization and clustering hyperparameters, and does not rely on finely tuned parameter choices.
+Table 9: Comparison of fusion strategies for multimodal integration based on TQ.
+     Model      Fusion Strategy      TQ
+     MultiGraph      Concatenation      0.225
+     MultiGraph      Attention      0.384
+     MultiGraph      Concatenation + MLP      0.165
+6.3 Random Seeds Stability and Fusion Strategy
+To evaluate robustness to stochastic initialization, we repeat the full baseline configuration using multiple random seeds.
+Table 8 reports the mean TQ, standard deviation, and 95% confidence interval.
+Both models exhibit low variance across runs, indicating that performance is stable and reproducible.
+We further compare attention-based fusion with alternative fusion strategies for the MultiGraph model, including plain concatenation and concatenation augmented with a multi-layer perceptron (MLP).
+As shown in Table 9, attention-based fusion yields substantially higher TQ than both concatenation-based variants, indicating that adaptive, context-aware weighting is more effective for integrating heterogeneous modalities than static or prior-driven fusion schemes.
+6.4 Architectural and graph construction choices
+In addition to the reported ablation and sensitivity analyses, we conducted preliminary validation experiments to assess the impact of architectural depth and graph construction choices on model behavior.
+In particular, we varied the number of GCN layers and attention heads within a small range (1–3) and observed no consistent or systematic changes in TQ across datasets.
+Based on this observation, these architectural parameters are fixed in all reported experiments in order to reduce the hyperparameter search space and improve reproducibility.
+We also explored alternative graph construction strategies, including global similarity and distance-based thresholding.
+These approaches were found to be sensitive to dataset scale and density, often requiring dataset-specific tuning to maintain stable sparsity levels.
+To avoid this dependency, we adopt a consistent top-$k$ neighborhood strategy for both semantic and geographic graphs throughout the paper.
+This choice yields predictable graph sparsity across datasets and allows performance differences to more directly reflect the proposed multimodal fusion mechanism and learning objectives rather than incidental graph construction effects.
+6.5 Overall Model Behavior
+Overall, the ablation, sensitivity, and stability analyses demonstrate that the proposed framework is robust across a wide range of configurations.
+Performance consistently peaks at the baseline configuration ($\alpha=0.8$, $\beta=0.2$, $\lambda=0.1$, $seed=42$, $k_{\text{means}}=15$, and $\tau=0.5$) while maintaining low variance across random seeds.
+These results indicate that the reported improvements are not artifacts of hyperparameter tuning or stochastic effects, but reflect stable and reproducible behavior of the proposed model.
 7 Case Study Evaluation
@@
 The dataset comprises 28,933 geo-located tweets posted within and around the area of Texas, obtained by filtering tweets based on Geographic Bounds defined by the polygon $\text{Polygon}(-98.77,27.58;-90.86,27.58;-90.86,30.66;-98.77,30.66;-98.77,27.58)$ and the Time Frame from August 25 to September 7, 2017.
+Figure 3: Topic-word distributions for a subset of the Hurricane Harvey dataset, applying MonoGraph (top) and the MultiGraph (bottom) models.
 7.1 Topic Interpretation
@@
 Weather-related discourse is further reflected in Topics 2, 7, 5 and 9, in MonoGraph and Topics 3, 4, 9, 6 and 8 in MultiGraph, featuring overlapping terms such as “pressure,” “gust,” “humidity,” “wind,” and “clouds.” While this lexical overlap reduces topic distinctiveness, it underscores the recurring presence of weather-related signals across the corpus, a common characteristic in disaster-related social media data.
-This also illustrates the trade-off of fixing the number of topics ($K=10$), which facilitates stability but can lead to partial redundancy.
+This also illustrates the trade-off of fixing the number of topics ($k_{\text{means}}=15$), which facilitates stability but can lead to partial redundancy.
 Finally, content-sharing and localized observational behavior emerge as a consistent pattern.
@@
 8 Discussion and Limitations
-Across all datasets, both MonoGraph and MultiGraph consistently outperformed baseline methods, achieving higher intra-cluster cohesion and lower inter-cluster similarity.
-Their relative strengths varied with event characteristics, suggesting sensitivity to spatial scale, discourse patterns, and data quality.
-Fixing the number of clusters at $K=10$ facilitated direct cross-model comparisons and system stability, but also introduced constraints.
-In some cases, semantically similar clusters emerged, particularly in weather-related topics, where recurring terms such as “pressure,” “humidity,” “wind” were dispersed across multiple clusters.
-This effect may be amplified in datasets containing unfiltered automated “weather bot” content.
-Empty or low-content clusters occasionally appeared due to sparse coverage, noise, or embedding collapse.
-The spatial distribution of events significantly shaped clustering performance.
-MultiGraph produced the lowest inter-cluster similarity for geographically concentrated events (Chile wildfires, Napa earthquake), benefiting from modality-specific graph encoding and cross-attention fusion.
-MonoGraph showed better separation in smaller or noisier AOIs (Ahr Valley floods, Hurricane Harvey), where its unified representation was more robust to diffuse or overlapping discourse.
-Spatially, MultiGraph generated compact, well-defined hotspots with fewer outliers, aligning clusters closely with localized disaster impacts, an asset for targeted response planning.
-MonoGraph captured broader spatial patterns, useful for representing large-scale narratives but with a tendency to overestimate geographic extent.
-Several limitations merit consideration.
-First, the framework’s performance is sensitive to the quality and balance of input modalities.
-Second, the current implementation does not explicitly model temporal dynamics, which constrains the ability to capture and analyze the evolution of topics over time.
-Third, the use of a fixed number of clusters ($K$) may not reflect the true topical complexity of diverse events, suggesting the need for adaptive or non-parametric clustering strategies.
-Finally, MultiGraph’s scalability to real-time, high-volume streams is untested.
-These constraints may hinder applicability to scenarios involving extremely high-frequency data or expansive geographic coverage unless complemented by distributed computing or streaming graph processing frameworks.
-Despite these constraints, both architectures effectively detect geographically specific and semantically coherent topics that align with real-world hazard footprints, offering practical value for enhancing situational awareness and supporting decision-making in disaster management.
+Across all datasets, both MonoGraph and MultiGraph consistently outperform baseline methods, achieving higher intra-cluster cohesion and lower inter-cluster similarity.
+Their relative strengths vary with event characteristics, reflecting sensitivity to spatial scale, discourse structure, and data quality.
+Fixing the number of clusters at $k_{\text{means}}=15$ facilitates direct cross-model comparison and stabilizes evaluation, but also introduces constraints.
+In particular, a fixed $k_{\text{means}}$ may not reflect the true topical complexity of heterogeneous events.
+In some cases, semantically similar clusters emerge, especially for weather-related topics where recurring terms such as ”pressure,” ”humidity” and ”wind” are distributed across multiple clusters.
+This effect may be further amplified in datasets containing unfiltered automated or bot-generated content.
+Occasional empty or low-content clusters also appear in sparsely populated or noisy regions, likely due to limited message coverage or embedding collapse.
+From a spatial perspective, event structure strongly shapes clustering behavior.
+MultiGraph produces compact, well-defined spatial clusters with fewer outliers in geographically concentrated events (e.g., Chile wildfires and the Napa earthquake), benefiting from modality-specific graph construction and cross-attention fusion.
+MonoGraph exhibits more robust separation in smaller or noisier areas of interest (e.g., Ahr Valley floods and Hurricane Harvey), where a unified representation better accommodates diffuse or overlapping discourse.
+As a result, MultiGraph tends to align clusters more closely with localized hazard impacts, while MonoGraph captures broader spatial narratives but may overestimate geographic extent.
+Several limitations of the proposed framework naturally motivate directions for future research.
+Model performance depends on the quality and balance of the input modalities, which may vary across platforms and events and affect generalization.
+Temporal dynamics are not explicitly modeled, limiting the ability to capture topic evolution and cascading effects over time.
+While fixing $k_{\text{means}}$ enables controlled comparison across methods, adaptive or stability-driven strategies for determining the number of clusters, such as consensus-based or non-parametric approaches, represent an important extension.
+In addition, automated or bot-generated content is not explicitly filtered, which may introduce redundancy or noise in certain datasets.
+Evaluation currently relies on intrinsic topic quality metrics and spatial statistics; incorporating human-centered assessment, such as expert judgment of topic interpretability and relevance, or downstream task-based validation, would further strengthen claims of practical utility.
+Finally, the scalability of the MultiGraph architecture to real-time, high-volume data streams remains untested and may require distributed or streaming graph-processing solutions, alongside exploration of alternative graph construction strategies, additional modalities (e.g., sentiment or temporal signals), and explicit social network structures to enhance community-level insights.
 9 Conclusion
@@
 Although the current implementation focuses on text and location, it can be extended to additional modalities such as sentiment, temporal dynamics, or imagery, enabling richer, more context-aware analyses in domains including disaster management, urban planning, and public health surveillance.
-Future directions include incorporating time-aware modeling, improving scalability for real-time data streams, and leveraging social network structures to enhance community-level insights.
-By unifying multimodal signals in an end-to-end learning model, this work advances toward more holistic and interpretable models for understanding the complex information landscapes of social media.
+By unifying multimodal semantic and spatial signals within an end-to-end learning framework, this work takes a step toward more holistic and interpretable models for understanding the complex information landscapes of geo-social media.
+[…]
+Acknowledgements
+This research was funded by the Austrian Research Promotion Agency (FFG) through the project MOSAIK (grant number 926200) and by the European Union under the Horizon Europe Research and Innovation Actions programme (grant agreement No.
+101093003, HORIZON-CL4-2022-DATA-01-01).
+Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission.
+Neither the European Union nor the European Commission can be held responsible for them.